### v1

# Elementary commutator method for the Dirac equation with long-range perturbations

Abstract
We present direct and elementary commutator techniques for the Dirac equation with long-range electric and mass perturbations. The main results are absence of generalized eigenfunctions and locally uniform resolvent estimates, both in terms of the optimal Besov-type spaces. With an additional massless assumption, we also obtain an algebraic radiation condition of projection type. For their proofs, following the scheme of Ito–Skibsted, we adopt, along with various weight functions, the generator of radial translations as conjugate operator, and avoid any of advanced functional analysis, pseudodifferential calculus, or even reduction to the Schrödinger equation. The results of the paper would serve as a foundation for the stationary scattering theory of the Dirac operator.
1 Settings and results
1.1 Settings
1.1.1 Introduction
In this paper we investigate the Dirac equation in the $d$-dimensional configuration space $\mathbb{R}^{d}$ with $d\in\mathbb{N}=\{1,2,\ldots\}$. We write it in the form of evolution equation as
$\partial_{t}\psi=-\mathrm{i}H\psi$
with
$H=\alpha_{j}p_{j}+q,\quad p_{j}=-\mathrm{i}\partial_{j}\ \text{for }j=1,\dots,d.$
Throughout the paper we adopt Einstein’s summation convention however without tensorial superscripts. Here, for some $n\in\mathbb{N}$, $\psi$ is an unknown $\mathbb{C}^{n}$-valued function, and $q$ is a given $n\times n$ Hermitian matrix-valued function, including simultaneously an electromagnetic potential and a mass term. We assume that $q$ is time-independent and long-range as will be formulated below. In addition, $\alpha_{1},\dots,\alpha_{d}$ are $n\times n$ Hermitian matrices satisfying the anti-commutation relations
$\{\alpha_{j},\alpha_{k}\}=2\delta_{jk}I_{n}\ \ \text{for }j,k=1,\dots,d,$ (1.1)
where $I_{n}$ is the identity matrix, which is omitted in the following. It is well-known that such matrices exist if $n=2^{[d/2]}$, or if $n=2^{[(d+1)/2]}$ when $\beta$ satisfying (1.8) below is added. However, our arguments are independent of $n$ or of explicit representations of $\alpha_{1},\dots,\alpha_{d}$, and dependent only on the above anti-commutation relations.
The purpose of the paper is to introduce a new commutator scheme to the analysis of the Dirac operator, following a paper by Ito–Skibsted [16], which originally applied it to the Schrödinger operator. The scheme is quite elementary. The conjugate operator is simply the generator of radial translations, and the main tools are the product rule of derivatives and the Cauchy–Schwartz inequality, combined with exquisite choice of weight functions. We do not need energy cut-offs from functional analysis or pseudodifferential calculus, or even reduction to the Schrödinger operator. See the papers [14, 38, 8, 9, 4] for the Mourre theory with more technical conjugate operators. Then the difficulty of the paper is that the commutator is merely a first-order differential operator, for which we cannot generally expect a sign. However, we can extract (weighted) positivity from it, which is the most technical novelty of the paper. We remark that a commutator argument of this type may be called the virial theorem [20, 37, 25, 26, 21, 13] or the method of multiplier [6], however we keep to use the terminology “commutator”.
The first result of the paper is Rellich’s theorem, or absence of generalized eigenfunctions in an optimally weighted Besov-type space, called the Agmon-Hörmander space. Although that of $L^{2}$-eigenfunctions has long been discussed in various settings [20, 30, 37, 25, 26, 21, 6, 7, 13], only Vogelsang [31] seems to have considered it in a wider weighted space $L^{2}_{-1/2}(\mathbb{R}^{d})$ for $d=3$. Hence we might for the first time extend it to the optimally largest space, which properly contains $L^{2}_{-1/2}(\mathbb{R}^{d})$. Note that, to this end, the choice of our conjugate operator is essential, while the generator of dilations, adopted in the ordinary Mourre theory, cannot go beyond $L^{2}(\mathbb{R}^{d})$. The second result is the local LAP bounds, or locally uniform resolvent estimates as operators between certain Besov-type spaces, which we also call the Agmon-Hörmander spaces. We obtain them with a long-range electric and mass perturbation in a general dimension. This subject also has a long history [33, 31, 1, 12], and recent interest is more in high energy, threshold or global estimates [36, 28, 14, 3, 22, 32, 4, 11, 5]. Our result is still local, however adoption of the Agmon-Hörmander spaces would be new for the Dirac operator of such generality. Finally, with an additional massless assumption, we verify Sommerfeld’s uniqueness for a solution to the associated Helmholtz equation. For that we present two types of radiation conditions, the analytic and the algebraic ones. The former is a natural analogue from the Schrödinger theory, extending a result by Pladdy–Saitō–Umeda [27] to long-range perturbations in the massless case. On the other hand, as for the latter, we construct an explicit orthogonal projections onto the outgoing and the incoming subspaces of $\mathbb{C}^{n}$. This extends a part of results by Kravchenko–Castillo P. [23] and Marmolejo-Olea–Pérez-Esteva [24]. Let us remark that also Vogelsang [31] seems to have discussed it in an indirect manner.
Whereas these results have their own value, they would also serve as fundamental ingredients for the stationary scattering theory, as in the Schrödinger case [17, 18]. So far, while the time-dependent scattering theory for the Dirac operator has been intensively studied [35, 29, 14, 8, 9, 2], the stationary theory seems to have been less developed, and done mainly for the short-range perturbations [1, 34]. Though Gâtel–Yafaev [12] took stationary approach to the long-range case, their wave operators are formulated in a time-dependent manner. Hopefully, we could discuss elsewhere the purely stationary scattering theory for long-range perturbations based on the results of the paper.
1.1.2 Long-range perturbations
To state precise assumptions on $q$, let us introduce a modified radius and the associated radial derivative. Choose $\chi\in C^{\infty}(\mathbb{R})$ such that
$\chi(t)=\left\{\begin{array}[]{ll}1&\mbox{ for }t\leq 1,\\ 0&\mbox{ for }t\geq 2,\end{array}\right.\quad\chi^{\prime}\leq 0,$ (1.4)
and define $f\in C^{\infty}(\mathbb{R}^{d})$ as
$f(x)=\chi(|x|)+|x|(1-\chi(|x|))\ \ \text{for }x\in\mathbb{R}^{d}.$
Obviously, $f$ coincides with $r=|x|$ for $|x|\geq 2$, but the values for $|x|<2$ are modified so that $f$ is uniformly positive on the whole $\mathbb{R}^{d}$. Thus, in particular, we may use $f$ as a weight function replacing the standard one $\langle x\rangle=(1+x^{2})^{1/2}$. In addition, we define
$\partial_{f}=(\partial_{j}f)\partial_{j},$ (1.5)
which is non-singular and globally defined on $\mathbb{R}^{d}$. We denote the set of all $n\times n$ Hermitian matrices by $\mathbb{C}^{n\times n}_{\mathrm{Her}}$, and its norm, arbitrarily fixed one, by $|\cdot|$.
Assumption 1.1.
The perturbation $q$ splits as
$q=q_{0}+q_{1}+q_{2};\quad q_{0},q_{1}\in(C^{1}\cap L^{\infty})(\mathbb{R}^{d};\mathbb{C}^{n\times n}_{\mathrm{Her}}),\ \ q_{2}\in L^{\infty}(\mathbb{R}^{d};\mathbb{C}^{n\times n}_{\mathrm{Her}}),$
such that the identities
$\widetilde{q}_{0}:=\alpha_{1}q_{0}\alpha_{1}=\alpha_{2}q_{0}\alpha_{2}=\dots=\alpha_{d}q_{0}\alpha_{d}$ (1.6)
hold on $\mathbb{R}^{d}$, and that for some $\rho\in(0,1]$ and $C>0$
$|q_{1}|\leq Cf^{-(1+\rho)/2},\quad|\partial_{1}q_{0}|+\dots+|\partial_{d}q_{0}|+|\partial_{f}q_{1}|+|q_{2}|\leq Cf^{-1-\rho}$ (1.7)
hold on $\mathbb{R}^{d}$.
Remarks 1.2.
  1. 1.
    The long-range part $q_{0}$ may be considered as a sum of electric potential and mass possibly with non-zero limit at spatial infinity. To see this, consider the ordinary Dirac equation with electromagnetic potential $A=(V,-A_{1},\dots,-A_{d})$ and mass $m$:
    $\partial_{t}\psi=-\mathrm{i}(\alpha_{j}p_{j}+\beta m+V-\alpha_{j}A_{j})\psi,$
    where $\beta\in\mathbb{C}^{n\times n}_{\mathrm{Her}}$ satisfies the anti-commutation relations
    $\{\alpha_{j},\beta\}=0\ \ \text{for }j=1,\dots,d,\quad\beta^{2}=1.$ (1.8)
    Our model covers this case with $q=\beta m+V-\alpha_{j}A_{j}$, and $q_{0}=V+\beta m$ in fact satisfies (1.6) with $\widetilde{q}_{0}=V-\beta m$.
  2. 2.
    The identity (1.6) can be relaxed to admit short-range errors.
  3. 3.
    We can further add local singularities, as long as $H$ with domain $C^{\infty}_{\mathrm{c}}(\mathbb{R}^{d};\mathbb{C}^{n})$ is essentially self-adjoint on $\mathcal{H}=L^{2}(\mathbb{R}^{d};\mathbb{C}^{n})$, and the unique continuation property is available. Note that the essential self-adjointness is required only for justification of doing and undoing commutators, as extensions from $C^{\infty}_{\mathrm{c}}(\mathbb{R}^{d};\mathbb{C}^{n})$. Such justification is not discussed in this paper, and we refer to the previous work [16].
1.1.3 Agmon–Hörmander spaces
All of our main results are stated in terms of the Agmon–Hörmander spaces, or the Besov spaces associated with the multiplication operator $f$ on $\mathcal{H}$. See also a paper by Jensen–Perry [19]. Here let us recall them. Let $F(S)$ be the characteristic function of a given subset $S\subset\mathbb{R}^{d}$, and set
$F_{\nu}=F\bigl(\{x\in\mathbb{R}^{d};\ 2^{\nu}\leq f<2^{\nu+1}\}\bigr)\quad\text{for }\nu\in\mathbb{N}_{0}=\{0\}\cup\mathbb{N}.$ (1.9)
Then we define function spaces $\mathcal{B}$, $\mathcal{B}^{*}$ and $\mathcal{B}^{*}_{0}$ as
$\mathcal{B}$ $=\bigl\{\psi\in L^{2}_{\mathrm{loc}}(\mathbb{R}^{d};\mathbb{C}^{n});\ \|\psi\|_{\mathcal{B}}<\infty\bigr\},\quad\|\psi\|_{\mathcal{B}}=\sum_{\nu\in\mathbb{N}_{0}}2^{\nu/2}\|F_{\nu}\psi\|_{{\mathcal{H}}},$
$\mathcal{B}^{*}$ $=\bigl\{\psi\in L^{2}_{\mathrm{loc}}(\mathbb{R}^{d};\mathbb{C}^{n});\ \|\psi\|_{\mathcal{B}^{*}}<\infty\bigr\},\quad\|\psi\|_{\mathcal{B}^{*}}=\sup_{\nu\in\mathbb{N}_{0}}2^{-\nu/2}\|F_{\nu}\psi\|_{{\mathcal{H}}},$
$\mathcal{B}^{*}_{0}$ $=\Bigl\{\psi\in\mathcal{B}^{*};\ \lim_{\nu\to\infty}2^{-\nu/2}\|F_{\nu}\psi\|_{{\mathcal{H}}}=0\Bigr\},$
respectively. Note that, if we denote the weighted $L^{2}$ spaces by
$L_{s}^{2}=f^{-s}L^{2}(\mathbb{R}^{d};\mathbb{C}^{n})\ \ \text{for }s\in\mathbb{R},$
then for any $s>1/2$ we have strict inclusions
$L^{2}_{s}\subsetneq\mathcal{B}\subsetneq L^{2}_{1/2}\subsetneq\mathcal{H}\subsetneq L^{2}_{-1/2}\subsetneq\mathcal{B}^{*}_{0}\subsetneq\mathcal{B}^{*}\subsetneq L^{2}_{-s}.$
1.2 Main results
1.2.1 Rellich’s theorem
Our first main theorem is Rellich’s theorem. We set
$m_{+}$ $=\liminf_{R\to\infty}\{\lambda\in\mathbb{R};\ \lambda\geq q_{0}(x)\ \text{for all }|x|\geq R\},$
$m_{-}$ $=\limsup_{R\to\infty}\{\lambda\in\mathbb{R};\ \lambda\leq q_{0}(x)\ \text{for all }|x|\geq R\},$
where the inequalities involving $\lambda$ and $q_{0}(x)$ are those for matrices, or quadratic forms on $\mathbb{C}^{n}$.
Theorem 1.3.
Suppose Assumption 1.1. If $\phi\in\mathcal{B}^{*}_{0}$ and $\lambda\in\mathbb{R}\setminus[m_{-},m_{+}]$ satisfy
$(H-\lambda)\phi=0\ \ \text{in the distributional sense,}$ (1.10)
then $\phi=0$ as a function on $\mathbb{R}^{d}$. In particular, the self-adjoint realization of $H$ has no eigenvalues outside $m_{-},m_{+}]$, i.e., $\sigma_{\mathrm{pp}}(H)\setminus[m_{-},m_{+}]=\emptyset$.
Remarks 1.4.
  1. 1.
    The function space $\mathcal{B}^{*}_{0}$ is optimal in the sense that there actually exists a non-trivial solution to (1.10) in $\mathcal{B}^{*}$, see Remark 1.14 for the construction. The largest function space so far is $L^{2}_{-1/2}(\mathbb{R}^{d})$ for $d=3$ due to Vogelsang [31], as far as the authors are aware of.
  2. 2.
    As remarked in a useful survey by Ito–Yamada [15], if $q$ is smooth and purely electric, we can reduce it to the Schrödinger operator. In fact, apply $\alpha^{j}p_{j}$ to (1.10) and use the identity (1.10) itself, and then we can deduce
    $\bigl(p_{j}p_{j}+[\alpha_{j}p_{j},q]-(\lambda-q)^{2}\bigr)\phi=0.$ (1.11)
    However, in general case it is not straightforward how to control $\alpha_{j}p_{j},q]$ since $\alpha_{j}$ and $q$ do not commute. We shall not go through this expression, but directly study the identity (1.10).
1.2.2 LAP bounds
We next discuss the LAP (Limiting Absorption Principle) bounds for the resolvent
$R(z)=(H-z)^{-1}\in\mathcal{L}(\mathcal{H})\ \ \text{for }z\in\mathbb{C}\setminus\mathbb{R}.$
For any compact interval $I\subset\mathbb{R}\setminus[m_{-},m_{+}]$ let us denote
$I_{\pm}=\bigl\{z=\lambda+\mathrm{i}\Gamma\in\mathbb{C};\ \lambda\in I,\ 0<\pm\sigma\Gamma\leq 1\bigr\},$ (1.12)
respectively, where $\sigma=+$ if $I\subset(m_{+},\infty)$, and $\sigma=$- if $I\subset(-\infty,m_{-})$.
Theorem 1.5.
Suppose Assumption 1.1, and let $I\subset\mathbb{R}\setminus[m_{-},m_{+}]$ be a compact interval. Then there exists $C>0$ such that for any $\phi=R(z)\psi$ with $z\in I_{\pm}$ and $\psi\in\mathcal{B}$
$\|\phi\|_{\mathcal{B}^{*}}+\|p_{1}\phi\|_{\mathcal{B}^{*}}+\dots+\|p_{d}\phi\|_{\mathcal{B}^{*}}\leq C\|\psi\|_{\mathcal{B}}.$
In particular, the self-adjoint realization of $H$ has no singular continuous spectrum outside $m_{-},m_{+}]$, i.e., $\sigma_{\mathrm{sc}}(H)\setminus[m_{-},m_{+}]=\emptyset$.
Remark 1.6.
For the proof of the LAP bounds it seems to have been typical to somehow reduce it to the (free) Schrödinger operator. In fact, similarly to (1.11), we can deduce from the identity $(H-z)\phi=\psi$ that
$\bigl(p_{j}p_{j}+[\alpha_{j}p_{j},q]-(z-q)^{2}\bigr)\phi=(\alpha_{j}p_{j}-q+z)\psi.$
However, again, we provide a direct proof, employing an elementary commutator method very similar to that for Rellich’s theorem. Rellich’s theorem itself also plays an essential role in contradiction argument in the proof of the LAP bounds.
1.2.3 Radiation condition bounds and applications
For further arguments we require the following massless assumption.
Assumption 1.7.
In addition to Assumption 1.1, there exists $C^{\prime}\geq 0$ such that
$|(q_{0}+\widetilde{q}_{0})q_{1}-q_{1}(q_{0}+\widetilde{q}_{0})|\leq C^{\prime}f^{-1-\rho},\quad|q_{0}-\widetilde{q}_{0}|\leq C^{\prime}f^{-\rho}$
hold on $\mathbb{R}^{d}$.
Remark 1.8.
In terms of the ordinary model given in Remark 1.2, $q_{0}=V+\beta m$ satisfies the former bound as it is. However, the latter requires $m$ to be of order $f^{-\rho}$, and this is exactly the massless assumption at spatial infinity.
The following radiation condition bounds describe radial oscillatory behavior of the resolvent at spatial infinity. Using $\sigma$ defined right after (1.12), let us set
$p_{f}=(\partial_{j}f)p_{j},\quad\alpha_{f}=(\partial_{j}f)\alpha_{j},\quad\pi_{\pm}=\tfrac{1}{2}(1\pm\sigma\alpha_{f}),$ (1.13)
respectively, cf. (1.5). Due to (1.1), $\pi_{\pm}$ for $|x|\geq 2$ are orthogonal projections on $\mathbb{C}^{n}$. They represent the projections onto the subspaces of outgoing/incoming components, respectively, as can be seen in the following.
Theorem 1.9.
Suppose Assumption 1.7, let $I\subset\mathbb{R}\setminus[m_{-},m_{+}]$ be a compact interval, and let $\kappa\in(0,\rho/2)$. Then there exists $C>0$ such that for any $\phi=R(z)\psi$ with $z\in I_{\pm}$ and $\psi\in L^{2}_{1/2+\kappa}$
$\|\pi_{\mp}\phi\|_{L^{2}_{-1/2+\kappa}}+\|(p_{f}-\alpha_{f}(z-q_{0}))\phi\|_{L^{2}_{-1/2+\kappa}}\leq C\|\psi\|_{L^{2}_{1/2+\kappa}},$
respectively.
Remarks 1.10.
  1. 1.
    The first term on the left-hand side corresponds to the algebraic radiation condition due to Kravchenko–Castillo P. [23] and Marmolejo-Olea–Pérez-Esteva [24]. While they discussed an exterior problem without electromagnetic potentials, we do it for long-range perturbations without obstacles under massless assumption. It would not be difficult to extend our result to the exterior case, even with some unbounded obstacles, following the setting of Ito–Skibsted [16], but we shall not elaborate it in the present paper. On the other hand, the analytic radiation condition, like the one by Pladdy–Saitō–Umeda [27], can be obtained by combining the first and the second terms on the left-hand side, see Corollary 1.13.
  2. 2.
    The proof of Theorem 1.9 is more or less different from Ito–Skibsted [16].
Then let us present applications of Theorems 1.3, 1.5 and 1.9. The first one is the LAP, or existence of the limits of resolvent as $I_{\pm}\ni z\to\lambda\in I$. Denote by $\mathcal{L}(X,Y)$ the space of all bounded operators from a Banach space $X$ to another $Y$.
Corollary 1.11.
Suppose Assumption 1.7, let $I\subset\mathbb{R}\setminus[m_{-},m_{+}]$ be a compact interval, and let $s\in(1/2,(1+\rho)/2)$ and $\epsilon\in(0,s-1/2)$. Then there exists $C>0$ such that for any $z,w\in I_{+}$ or $z,w\in I_{-}$
$\begin{split}\|R(z)-R(w)\|_{\mathcal{L}(L^{2}_{s},L^{2}_{-s})}+\sum_{j=1}^{d}\|p_{j}R(z)-p_{j}R(w)\|_{\mathcal{L}(L^{2}_{s},L^{2}_{-s})}\leq C|z-w|^{\epsilon}.\end{split}$ (1.14)
In particular, $R(z),p_{1}R(z),\dots,p_{d}R(z)$ have uniform limits as $I_{\pm}\ni z\to\lambda\in I$ in the norm topology of $\mathcal{L}(L^{2}_{s},L^{2}_{-s})$. In addition, if one denotes
$\begin{split}R_{\pm}(\lambda)=\lim_{I_{\pm}\ni z\to\lambda}R(z)\ \ \text{in }\mathcal{L}(L^{2}_{s},L^{2}_{-s}),\end{split}$ (1.15)
respectively, then $R_{\pm}(\lambda),p_{1}R_{\pm}(\lambda),\dots,p_{d}R_{\pm}(\lambda)$ belong to $\mathcal{L}(\mathcal{B},\mathcal{B}^{*})$.
Theorem 1.9 immediately extends to $R_{\pm}(\lambda)$ as follows.
Corollary 1.12.
Suppose Assumption 1.7, let $I\subset\mathbb{R}\setminus[m_{-},m_{+}]$ be a compact interval, and let $\kappa\in(0,\rho/2)$. Then there exists $C>0$ such that for any $\phi=R_{\pm}(\lambda)\psi$ with $\lambda\in I$ and $\psi\in L^{2}_{1/2+\kappa}$
$\|\pi_{\mp}\phi\|_{L^{2}_{-1/2+\kappa}}+\|(p_{f}-\alpha_{f}(\lambda-q_{0}))\phi\|_{L^{2}_{-1/2+\kappa}}\leq C\|\psi\|_{L^{2}_{1/2+\kappa}},$
respectively.
The following Sommerfeld’s uniqueness characterizes $R_{\pm}(\lambda)$ as solution operators to the Helmholtz equation with outgoing or incoming radiation conditions.
Corollary 1.13.
Suppose Assumption 1.7, and let $\lambda\in\mathbb{R}\setminus[m_{-},m_{+}]$ and $\kappa\in(0,\rho/2)$. If $\psi\in L^{2}_{1/2+\kappa}$, then $\phi:=R_{\pm}(\lambda)\psi\in\mathcal{B}^{*}$, and it satisfies
  1. 1.
    $(H-\lambda)\phi=\psi$ in the distributional sense,
  2. 2.
    $\pi_{\mp}\phi\in L^{2}_{-1/2+\kappa}$, and $(p_{f}\mp\sigma(\lambda-q_{0}))\phi\in L^{2}_{-1/2+\kappa}$,
respectively. Conversely, if $\psi\in f^{-\kappa}\mathcal{B}$ and $\phi\in f^{\kappa}\mathcal{B}^{*}$ satisfy the above 1 and
  1. 3.
    $\pi_{\mp}\phi\in f^{-\kappa}\mathcal{B}_{0}^{*}$, or $(p_{f}\mp\sigma(\lambda-q_{0}))\phi\in f^{-\kappa}\mathcal{B}_{0}^{*}$,
then $\phi=R_{\pm}(\lambda)\psi$, respectively.
Remark 1.14.
Using Corollary 1.13, we can construct a non-trivial solution to (1.10). In fact, for any $\psi\in L^{2}_{1/2+\kappa}$ with $\kappa\in(0,\rho/2)$ we set
$\phi=R_{+}(\lambda)\psi-R_{-}(\lambda)\psi.$
Then it satisfies (1.10) due to 1 of Corollary 1.13, and is non-trivial due to 2 of Corollary 1.13 and Theorem 1.3.
2 Proof of Rellich’s theorem
2.1 Reduction to three propositions
In this section we prove Theorem 1.3, following the scheme of Ito–Skibsted [16]. We split the proof into three steps. Obviously, Theorem 1.3 is a consequence of the following propositions. Throughout the section we assume Assumption 1.1.
Proposition 2.1.
Let $\phi\in\mathcal{B}^{*}_{0}$ and $\lambda\in\mathbb{R}\setminus[m_{-},m_{+}]$ satisfy
$(H-\lambda)\phi=0\ \ \text{in the distributional sense.}$
Then there exists $\mu>0$ such that $\mathrm{e}^{\mu f}\phi\in\mathcal{B}_{0}^{*}$.
Proposition 2.2.
Let $\phi\in\mathcal{B}_{0}^{*}$ and $\lambda\in\mathbb{R}\setminus[m_{-},m_{+}]$ satisfy
  1. 1.
    $(H-\lambda)\phi=0$ in the distributional sense,
  2. 2.
    $\mathrm{e}^{\kappa f}\phi\in\mathcal{B}_{0}^{*}$ for some $\kappa>0$.
Then for any $\mu>0$ one has $\mathrm{e}^{(\kappa+\mu)f}\phi\in\mathcal{B}^{*}_{0}$.
Proposition 2.3.
Let $\phi\in\mathcal{B}_{0}^{*}$ and $\lambda\in\mathbb{R}\setminus[m_{-},m_{+}]$ satisfy
  1. 1.
    $(H-\lambda)\phi=0$ in the distributional sense,
  2. 2.
    $\mathrm{e}^{\kappa f}\phi\in\mathcal{B}_{0}^{*}$ for any $\kappa\geq 0$.
Then $\phi=0$ as a function on $\mathbb{R}^{d}$.
We will prove these propositions in Sections 2.3, 2.4 and 2.5, respectively, after short preliminaries in Section $2.$
2.2 Preliminaries
2.2.1 Conjugate operator
Here we present formulas, to be repeatedly referred to in this and later sections, involving a general weight function. The proofs of Propositions 2.1, 2.2 and 2.3 depend on a commutator method sharing conjugate operators of the same form, however, with different weight functions.
Given a weight function $\Theta$, we introduce a radial differential operator
$A_{\Theta}=2\mathop{\mathrm{Re}}(\Theta p_{f})=\Theta p_{f}+p_{f}^{*}\Theta,$ (2.1)
where $p_{f}^{*}$ denotes the adjoint of $p_{f}$ from (1.13), unlike the physical convention. The function $\Theta$ will be specified soon below, but for the moment we only assume the following properties. We let $\Theta$ be a smooth function only of $f$ satisfying
$\Theta\geq 0,\quad\mathop{\mathrm{supp}}\Theta\subset\{x\in\mathbb{R}^{d};\ |x|\geq 2\},\quad|\Theta^{(k)}|\leq C_{k}\ \ \text{for any }k\in\mathbb{N}_{0},$ (2.2)
with $\Theta^{(k)}$ being the $k$-th derivative of $\Theta$ in $f$. Due to the supporting property of $\Theta$, as far as $\Theta$ is involved, we may always identify $f$ with $|x|$.
Recalling notation (1.13), we will often use a decomposition
$\alpha_{j}p_{j}=\alpha_{f}p_{f}+\alpha_{j}\ell_{jk}p_{k};\quad\ell_{jk}=\delta_{jk}-(\partial_{j}f)(\partial_{k}f).$ (2.3)
In particular, we note that on $\mathop{\mathrm{supp}}\Theta$ the matrix $(\ell_{ij})_{i,j=1,\dots,d}$ represents the orthogonal projection onto the spherical direction, and moreover
$\ell_{jk}=f(\partial_{j}\partial_{k}f)\ \ \text{for }j,k=1,\dots,d.$ (2.4)
Now let us compute a commutator $2\mathrm{i}[H,A_{\Theta}]=\mathop{\mathrm{Im}}(A_{\Theta}H)$ for general $\Theta$.
Lemma 2.4.
One has the identity
$\mathop{\mathrm{Im}}(A_{\Theta}H)$ $=\mathop{\mathrm{Re}}(\alpha_{f}\Theta^{\prime}p_{f})+\mathop{\mathrm{Re}}(\alpha_{j}f^{-1}\Theta\ell_{jk}p_{k}\bigr)$
$\phantom{{}={}}{}-(\partial_{f}(q_{0}+q_{1}))\Theta+(\Delta f)q_{2}\Theta+q_{2}\Theta^{\prime}-2\mathop{\mathrm{Im}}(q_{2}\Theta p_{f}).$
Remark 2.5.
In most of the later applications we will squeeze (weighted) positivity from the first and second terms on the right-hand side without energy cut-offs, in spite that they are first order differential operators with different weights $\Theta^{\prime}$ and $f^{-1}\Theta$. This is the most technical novelty of the paper. The other terms along with the left-hand side are negligible. For instance, the expectation of $\mathop{\mathrm{Im}}(A_{\Theta}H)=\mathop{\mathrm{Im}}(A_{\Theta}(H-\lambda))$ vanishes on an eigenstate $\phi$.
Proof.
We decompose the left-hand side of the asserted identity as
$\mathop{\mathrm{Im}}(A_{\Theta}H)=\mathop{\mathrm{Im}}(A_{\Theta}\alpha_{j}p_{j})+\mathop{\mathrm{Im}}(A_{\Theta}(q_{0}+q_{1}))+\mathop{\mathrm{Im}}(A_{\Theta}q_{2}).$ (2.5)
We can compute the first term on the right-hand side of (2.5) by using (2.4) as
$\mathop{\mathrm{Im}}(A_{\Theta}\alpha_{j}p_{j})$ $=\mathop{\mathrm{Im}}\bigl((\Theta p_{f}+p_{f}^{*}\Theta)\alpha_{j}p_{j}\bigr)$
$=\mathop{\mathrm{Im}}(\alpha_{j}\Theta p_{f}p_{j})-\mathop{\mathrm{Im}}(\alpha_{j}p_{j}\Theta p_{f})$
$=\mathop{\mathrm{Re}}(\alpha_{f}\Theta^{\prime}p_{f})+\mathop{\mathrm{Re}}\bigl(\alpha_{j}\Theta(\partial_{j}\partial_{k}f)p_{k}\bigr)$
$=\mathop{\mathrm{Re}}(\alpha_{f}\Theta^{\prime}p_{f})+\mathop{\mathrm{Re}}(\alpha_{j}f^{-1}\Theta\ell_{jk}p_{k}\bigr).$
As for the second and third terms on the right-hand side of (2.5), it is easy to see that
$\mathop{\mathrm{Im}}(A_{\Theta}(q_{0}+q_{1}))=-(\partial_{f}(q_{0}+q_{1}))\Theta,$
and that
$\mathop{\mathrm{Im}}(A_{\Theta}q_{2})$ $=\mathop{\mathrm{Im}}\bigl((\Theta p_{f}+p_{f}^{*}\Theta)q_{2}\bigr)$
$=\mathop{\mathrm{Re}}\bigl((\partial_{j}\Theta\partial_{j}f)q_{2}\bigr)+2\mathop{\mathrm{Im}}(p_{f}^{*}\Theta q_{2})$
$=(\Delta f)q_{2}\Theta+q_{2}\Theta^{\prime}-2\mathop{\mathrm{Im}}(q_{2}\Theta p_{f}).$
Hence we obtain the assertion. ∎
We will also often use the following identity.
Lemma 2.6.
For any $z\in\mathbb{C}$ and $\Xi\in C^{2}(\mathbb{R}^{d};\mathbb{R})$ one can write
$p_{j}\Xi p_{j}$ $=(z-q)^{*}(z-q)\Xi-\mathop{\mathrm{Im}}\bigl(\alpha_{j}(\partial_{j}\Xi)(z-q)\bigr)+\tfrac{1}{2}(\Delta\Xi)$
$\phantom{={}}{}+\mathop{\mathrm{Re}}\bigl(\bigl(2(z-q)^{*}\Xi+\mathrm{i}\alpha_{j}(\partial_{j}\Xi)\bigr)(H-z)\bigr)+(H-z)^{*}\Xi(H-z).$
Remark 2.7.
This implies that $p_{j}\Xi p_{j}$ is comparable to $(z-q)^{*}(z-q)\Xi\sim\Xi$. In fact, the second and the third term are of lower order since $\Xi$ is differentiated, and the third and the fourth are negligible due to the factor (Hz)(H-z).
Proof.
We can directly compute it as follows. In fact, by using (1.1) we have
$p_{j}\Xi p_{j}$ $=\mathop{\mathrm{Re}}(\Xi p_{j}p_{j})+\mathop{\mathrm{Im}}\bigl((\partial_{j}\Xi)p_{j}\bigr)$
$=\mathop{\mathrm{Re}}(\alpha_{j}\alpha_{k}\Xi p_{j}p_{k})+\tfrac{1}{2}(\Delta\Xi)$
$=p_{j}\alpha_{j}\Xi\alpha_{k}p_{k}-\mathop{\mathrm{Im}}\bigl(\alpha_{j}(\partial_{j}\Xi)\alpha_{k}p_{k}\bigr)+\tfrac{1}{2}(\Delta\Xi)$
$=(z-q+H-z)^{*}\Xi(z-q+H-z)$
$\phantom{={}}{}-\mathop{\mathrm{Im}}\bigl(\alpha_{j}(\partial_{j}\Xi)(z-q+H-z)\bigr)+\tfrac{1}{2}(\Delta\Xi)$
$=(z-q)^{*}(z-q)\Xi+2\mathop{\mathrm{Re}}\bigl((z-q)^{*}\Xi(H-z)\bigr)+(H-z)^{*}\Xi(H-z)$
$\phantom{={}}{}-\mathop{\mathrm{Im}}\bigl(\alpha_{j}(\partial_{j}\Xi)(z-q)\bigr)-\mathop{\mathrm{Im}}\bigl(\alpha_{j}(\partial_{j}\Xi)(H-z)\bigr)+\tfrac{1}{2}(\Delta\Xi).$
Thus obtain the assertion. ∎
2.2.2 Weight functions for Rellich’s theorem
The weight function we use in this section is of the form
$\Theta=\Theta_{a,b,\nu}^{\kappa,\mu}=\chi_{a,b}\mathrm{e}^{\theta},$ (2.6)
which is dependent on parameters $a,b,\nu\in\mathbb{N}_{0}$ and $\kappa,\mu\geq 0$. Here $\chi_{a,b}\in C^{\infty}(\mathbb{R}^{d})$ is a smooth cutoff function to a dyadic annular domain, defined as
$\chi_{a,b}=\bar{\chi}_{a}\chi_{b};\quad\bar{\chi}_{a}=1-\chi(f/2^{a}),\quad\chi_{b}=\chi(f/2^{b}),$ (2.7)
with $\chi\in C^{\infty}(\mathbb{R})$ being from (1.4). The exponent $\theta\in C^{\infty}(\mathbb{R}^{d})$ is defined as
$\theta=\theta_{\nu}^{\kappa,\mu}=2\kappa f+2\mu\int_{0}^{f}(1+\tau/2^{\nu})^{-1-\delta}\,\mathrm{d}\tau.$ (2.8)
Here and below the constant $\delta\in(0,\rho)$ is arbitrarily fixed, so that the dependence on it is suppressed in (2.6). Note that the integral in (2.8) is a refinement of the Yosida approximation, satisfying
$0<\int_{0}^{f}(1+\tau/2^{\nu})^{-1-\delta}\,\mathrm{d}\tau\leq\min\{f,\delta 2^{\nu}\}\ \ \text{for any }\nu\in\mathbb{N}_{0},$
and
$\int_{0}^{f}(1+\tau/2^{\nu})^{-1-\delta}\,\mathrm{d}\tau\uparrow f\ \ \text{pointwise as }\nu\to\infty.$
We note that, if we denote the derivatives in $f$ by primes as before, then
$\begin{split}\theta^{\prime}=2\kappa+2\mu(1+f/2^{\nu})^{-1-\delta},\quad\theta^{\prime\prime}=-2(1+\delta)\mu 2^{-\nu}(1+f/2^{\nu})^{-2-\delta},\quad\ldots.\end{split}$
In particular, noting $2^{-\nu}(1+f/2^{\nu})^{-1}\leq f^{-1}$, we can bound for any $k\in\mathbb{N}$
$|(\theta-2\kappa f)^{(k)}|\leq C_{\delta,k}f^{-k}(\theta^{\prime}-2\kappa).$
2.3 A priori exponential decay estimate
Now we present a key lemma for the proof of Proposition 2.1.
Lemma 2.8.
Let $\lambda\gtrless m_{\pm}$, respectively, and let $\kappa=0$ in the definition (2.6) of $\Theta$. Then there exist $\mu,c,C>0$ and $a_{0}\in\mathbb{N}_{0}$ such that uniformly in $b>a\geq a_{0}$ and $\nu\in\mathbb{N}_{0}$
$\pm\mathop{\mathrm{Im}}\bigl(A_{\Theta}(H-\lambda)\bigr)$ $\geq c\min\{f^{-1},\theta^{\prime}\}\Theta-C\bigl(\chi_{a-1,a+1}^{2}+\chi_{b-1,b+1}^{2}\big)f^{-1}\mathrm{e}^{\theta}$
$\phantom{={}}{}-(H-\lambda)\gamma(H-\lambda),$
respectively, where $\gamma=\gamma_{a,b}$ is a certain function satisfying $\mathop{\mathrm{supp}}\gamma\subset\mathop{\mathrm{supp}}\chi_{a,b}$ and $|\gamma|\leq C_{a,b}$.
Remarks 2.9.
  1. 1.
    We can see that the first term on the right-hand side is exactly from the first and second term of Lemma 2.4 with $\Theta$ of the form (2.6).
  2. 2.
    Let us outline the deduction of Proposition 2.1 from Lemma 2.8: Take the expectation of the above inequality on the eigenstate $\phi$, and the left-hand side and the last term on the right-hand side vanish; Let $b\to\infty$, and a contribution from $\chi_{b-1,b+1}^{2}f^{-1}\mathrm{e}^{\theta}$ vanishes by definition of $\mathcal{B}^{*}_{0}$; Then let $\nu\to\infty$, which replaces the Yosida approximation $\theta$ by $f$. Proposition 2.1 thus follows.
Proof.
Let us only discuss the upper sign since the lower one can be done by the same manner. Fix $\lambda>m_{\pm}$ and $\kappa=0$ as in the assertion. For the moment we merely choose $a_{0}\in\mathbb{N}_{0}$ such that for some $c_{1}>0$
$\lambda-q_{0}\geq c_{1}\ \ \text{uniformly in }\{x\in\mathbb{R}^{d};\ |x|\geq 2^{a_{0}}\}.$ (2.9)
Then, except for the last part of the proof, the below estimates are all uniform in $\mu\in(0,1]$, $b>a\geq a_{0}$ and $\nu\in\mathbb{N}_{0}$, and thus until then $c_{*},C_{*}>0$ are independent of them. Only in the very last step we shall restrict ranges of all these parameters to verify the assertion, where we may also retake $a_{0}\in\mathbb{N}_{0}$ larger if necessary.
In this proof we gather and absorb admissible error terms into
$\begin{split}Q&=\bigl(\chi_{a,b}f^{-1}\theta^{\prime}+\chi_{a,b}f^{-1-\rho}+|\chi_{a,b}^{\prime}|\bigr)\mathrm{e}^{\theta}\\ &\phantom{{}={}}{}+p_{j}\bigl(\chi_{a,b}f^{-1}\theta^{\prime}+\chi_{a,b}f^{-1-\rho}+|\chi_{a,b}^{\prime}|\bigr)\mathrm{e}^{\theta}p_{j}+(H-\lambda)f^{1+\rho}\Theta(H-\lambda),\end{split}$ (2.10)
which we prove at last is negligible.
Now let us start to compute the left-hand side of the asserted inequality. By Lemma 2.4 and the Cauchy–Schwarz inequality we first have
$\begin{split}\mathop{\mathrm{Im}}\bigl(A_{\Theta}(H-\lambda)\bigr)&=\mathop{\mathrm{Re}}\bigl(\alpha_{f}\theta^{\prime}\Theta p_{f}\bigr)+\mathop{\mathrm{Re}}\bigl(\alpha_{f}\chi_{a,b}^{\prime}\mathrm{e}^{\theta}p_{f}\bigr)+\mathop{\mathrm{Re}}\bigl(\alpha_{j}f^{-1}\Theta\ell_{jk}p_{k}\bigr)\\ &\phantom{{}={}}{}-\bigl(\partial_{f}(q_{0}+q_{1})\bigr)\Theta+(\Delta f)q_{2}\Theta+q_{2}\Theta^{\prime}-2\mathop{\mathrm{Im}}(q_{2}\Theta p_{f})\\ &\geq\mathop{\mathrm{Re}}\bigl(\alpha_{f}\theta^{\prime}\Theta p_{f}\bigr)+\mathop{\mathrm{Re}}\bigl(\alpha_{j}f^{-1}\Theta\ell_{jk}p_{k}\bigr)-C_{1}Q.\end{split}$ (2.11)
Recalling the expression (2.3) of $\ell_{jk}$, and using the Cauchy–Schwarz inequality and (2.9), we can combine and bound the first and second terms on the right-hand side of (2.11) as
$\begin{split}&\mathop{\mathrm{Re}}\bigl(\alpha_{f}\theta^{\prime}\Theta p_{f}\bigr)+\mathop{\mathrm{Re}}\bigl(\alpha_{j}f^{-1}\Theta\ell_{jk}p_{k}\bigr)\\ &=\tfrac{1}{2}\mathop{\mathrm{Re}}\bigl(\alpha_{f}(\theta^{\prime}+f^{-1})\Theta p_{f}\bigr)-\tfrac{1}{2}\mathop{\mathrm{Re}}\bigl(\alpha_{f}(f^{-1}-\theta^{\prime})\Theta p_{f}\bigr)\\ &\phantom{{}={}}+\tfrac{1}{2}\mathop{\mathrm{Re}}\bigl(\alpha_{j}(\theta^{\prime}+f^{-1})\Theta\ell_{jk}p_{k}\bigr)+\tfrac{1}{2}\mathop{\mathrm{Re}}\bigl(\alpha_{j}(f^{-1}-\theta^{\prime})\Theta\ell_{jk}p_{k}\bigr)\\ &=\tfrac{1}{2}\mathop{\mathrm{Re}}\bigl(\alpha_{j}(\theta^{\prime}+f^{-1})\Theta p_{j}\bigr)-\tfrac{1}{2}\mathop{\mathrm{Re}}\bigl(\alpha_{j}(f^{-1}-\theta^{\prime})\Theta h_{jk}p_{k}\bigr)\\ &=\tfrac{1}{2}(\lambda-q)(\theta^{\prime}+f^{-1})\Theta+\tfrac{1}{2}\mathop{\mathrm{Re}}\bigl((\theta^{\prime}+f^{-1})\Theta(H-\lambda)\bigr)\\ &\phantom{{}={}}-\tfrac{1}{2}\mathop{\mathrm{Re}}\bigl(\mathrm{e}^{\theta/2}\chi_{a,b}(f^{-1}-\theta^{\prime})\alpha_{j}h_{jk}p_{k}\mathrm{e}^{\theta/2}\bigr)\\ &\geq\tfrac{1}{2}(\lambda-q_{0}-q_{1})(\theta^{\prime}+f^{-1})\Theta-\tfrac{1}{4}(\lambda-q_{0})|f^{-1}-\theta^{\prime}|\Theta\\ &\phantom{{}={}}-\tfrac{1}{4}\mathrm{e}^{\theta/2}p_{i}h_{il}\alpha_{l}\chi_{a,b}(\lambda-q_{0})^{-1}|f^{-1}-\theta^{\prime}|\alpha_{j}h_{jk}p_{k}\mathrm{e}^{\theta/2}-C_{2}Q,\end{split}$ (2.12)
where we have set
$h_{jk}=(\partial_{j}f)(\partial_{k}f)-\ell_{jk}.$ (2.13)
We are going to further compute and bound the third term on the right-hand side of (2.12), which requires a subtle treatment. Let us set for short
$\Xi$ $=\chi_{a,b}(\lambda-q_{0})^{-1}\bigl((f^{-1}-\theta^{\prime})^{2}+f^{-2-2\rho}\bigr)^{1/2},$
$\widetilde{\Xi}$ $=\chi_{a,b}(\lambda-\widetilde{q}_{0})^{-1}\bigl((f^{-1}-\theta^{\prime})^{2}+f^{-2-2\rho}\bigr)^{1/2},$
cf. (1.6), and we proceed with the third term of (2.12) by using (1.1), (2.3) and the Cauchy–Schwarz inequality as
$\begin{split}&-\tfrac{1}{4}\mathrm{e}^{\theta/2}p_{i}h_{il}\alpha_{l}\chi_{a,b}(\lambda-q_{0})^{-1}|f^{-1}-\theta^{\prime}|\alpha_{j}h_{jk}p_{k}\mathrm{e}^{\theta/2}\\ &\geq-\tfrac{1}{4}\mathrm{e}^{\theta/2}p_{i}h_{il}\alpha_{l}\Xi\alpha_{j}h_{jk}p_{k}\mathrm{e}^{\theta/2}\\ &=-\tfrac{1}{4}\mathop{\mathrm{Re}}\bigl(\alpha_{l}\alpha_{j}\widetilde{\Xi}\mathrm{e}^{\theta/2}h_{il}h_{jk}p_{i}p_{k}\mathrm{e}^{\theta/2}\bigr)-\tfrac{1}{4}\mathop{\mathrm{Im}}\bigl(\alpha_{l}\mathrm{e}^{\theta/2}(\partial_{i}h_{il}\Xi h_{jk})\alpha_{j}p_{k}\mathrm{e}^{\theta/2}\bigr)\\ &=-\tfrac{1}{4}\mathop{\mathrm{Re}}\bigl(\widetilde{\Xi}\mathrm{e}^{\theta/2}h_{ij}h_{jk}p_{i}p_{k}\mathrm{e}^{\theta/2}\bigr)-\tfrac{1}{4}\mathop{\mathrm{Im}}\bigl(\alpha_{l}(\partial_{i}h_{il}\Xi h_{jk})\alpha_{j}\mathrm{e}^{\theta}p_{k}\bigr)\\ &\phantom{{}={}}+\tfrac{1}{8}\mathop{\mathrm{Re}}\bigl(\alpha_{l}(\partial_{i}h_{il}\Xi h_{jk})\alpha_{j}(\partial_{k}f)\theta^{\prime}\mathrm{e}^{\theta}\bigr)\\ &\geq-\tfrac{1}{4}\mathop{\mathrm{Re}}\bigl(\widetilde{\Xi}\mathrm{e}^{\theta/2}\delta_{ik}p_{i}p_{k}\mathrm{e}^{\theta/2}\bigr)-C_{3}Q.\end{split}$ (2.14)
Using (1.1) and the Cauchy–Schwarz inequality again, and noting $\alpha_{f}^{2}=1$, we continue (2.14) as
$\begin{split}&-\tfrac{1}{4}\mathrm{e}^{\theta/2}p_{i}h_{il}\alpha_{l}\chi_{a,b}(\lambda-q_{0})^{-1}|f^{-1}-\theta^{\prime}|\alpha_{j}h_{jk}p_{k}\mathrm{e}^{\theta/2}\\ &\geq-\tfrac{1}{4}\mathop{\mathrm{Re}}\bigl(\widetilde{\Xi}\mathrm{e}^{\theta/2}\alpha_{i}\alpha_{k}p_{i}p_{k}\mathrm{e}^{\theta/2}\bigr)-C_{3}Q\\ &=-\tfrac{1}{4}\mathrm{e}^{\theta/2}p_{i}\alpha_{i}\Xi\alpha_{k}p_{k}\mathrm{e}^{\theta/2}+\tfrac{1}{4}\mathop{\mathrm{Im}}\bigl(\mathrm{e}^{\theta/2}\alpha_{i}(\partial_{i}\Xi)\alpha_{k}p_{k}\mathrm{e}^{\theta/2}\bigr)-C_{3}Q\\ &=-\tfrac{1}{4}\bigl(p_{i}\alpha_{i}+\tfrac{\mathrm{i}}{2}\alpha_{f}\theta^{\prime}\bigr)\Xi\mathrm{e}^{\theta}\bigl(\alpha_{k}p_{k}-\tfrac{\mathrm{i}}{2}\alpha_{f}\theta^{\prime}\bigr)+\tfrac{1}{4}\mathop{\mathrm{Im}}\bigl(\alpha_{i}(\partial_{i}\Xi)\mathrm{e}^{\theta}\alpha_{k}p_{k}\bigr)\\ &\phantom{{}={}}-\tfrac{1}{8}\mathop{\mathrm{Re}}\bigl(\alpha_{i}\alpha_{f}(\partial_{i}\Xi)\theta^{\prime}\mathrm{e}^{\theta}\bigr)-C_{3}Q\\ &\geq-\tfrac{1}{4}p_{i}\alpha_{i}\Xi\mathrm{e}^{\theta}\alpha_{k}p_{k}+\tfrac{1}{8}\alpha_{i}\alpha_{k}\bigl(\partial_{i}\Xi\theta^{\prime}(\partial_{k}f)\mathrm{e}^{\theta}\bigr)-\tfrac{1}{16}\Xi\theta^{\prime 2}\mathrm{e}^{\theta}-C_{4}Q\\ &\geq-\tfrac{1}{4}(\lambda-q+H-\lambda)\Xi\mathrm{e}^{\theta}(\lambda-q+H-\lambda)+\tfrac{1}{16}\Xi\theta^{\prime 2}\mathrm{e}^{\theta}-C_{5}Q\\ &\geq-\tfrac{1}{4}(\lambda-q)\Xi\mathrm{e}^{\theta}(\lambda-q)-C_{6}Q\\ &\geq-\tfrac{1}{4}(\lambda-q_{0}-2q_{1})|f^{-1}-\theta^{\prime}|\Theta-C_{7}Q.\end{split}$ (2.15)
Hence by (2.11), (2.12) and (2.15) we obtain
$\begin{split}\mathop{\mathrm{Im}}\bigl(A_{\Theta}(H-\lambda)\bigr)&\geq(c_{1}-q_{1})\min\{f^{-1},\theta^{\prime}\}\Theta-C_{8}Q.\end{split}$ (2.16)
Next we discuss $Q$. Here we claim
$\begin{split}Q&\leq C_{9}\bigl(f^{-1}\theta^{\prime}+f^{-1-\rho}\bigr)\Theta+C_{9}\bigl(\chi_{a-1,a+1}^{2}+\chi_{b-1,b+1}^{2}\big)f^{-1}\mathrm{e}^{\theta}\\ &\phantom{{}={}}{}+C_{9}(H-\lambda)f^{1+\rho}\Theta(H-\lambda),\end{split}$ (2.17)
and for that it suffices to consider the second term of (2.10) since the contribution from $|\chi_{a,b}^{\prime}|\mathrm{e}^{\theta}$ is obviously absorbed into the second term of (2.17). However, the second term of (2.10) is obviously absorbed into the right-hand side of (2.17) due to Lemma 2.6 and the Cauchy–Schwarz inequality. Thus we obtain (2.17).
Therefore by (2.16) and (2.17) it follows that
$\begin{split}\mathop{\mathrm{Im}}\bigl(A_{\Theta}(H-\lambda)\bigr)&\geq\bigl((c_{1}-q_{1})\min\{f^{-1},\theta^{\prime}\}-C_{10}f^{-1}\theta^{\prime}-C_{10}f^{-1-\rho}\bigr)\Theta\\ &\phantom{{}={}}{}-C_{10}\bigl(\chi_{a-1,a+1}^{2}+\chi_{b-1,b+1}^{2}\big)f^{-1}\mathrm{e}^{\theta}\\ &\phantom{{}={}}{}-C_{10}(H-\lambda)f^{1+\rho}\Theta(H-\lambda),\end{split}$
and we finally discuss the coefficient of the first term on the right-hand side of (2.25). Now we restrict the ranges of parameters. We first retake $a_{0}\in\mathbb{N}_{0}$ larger if necessary, so that for some $c_{2}>0$
$(c_{1}-q_{1})\geq 2c_{2}\ \ \text{uniformly for }|x|\geq 2^{a_{0}},$
or on $\mathop{\mathrm{supp}}\Theta=\mathop{\mathrm{supp}}\chi_{a,b}$
$(c_{1}-q_{1})\min\{f^{-1},\theta^{\prime}\}-C_{10}f^{-1}\theta^{\prime}-C_{10}f^{-1-\rho}$
$\geq\min\bigl\{(c_{2}-C_{10}\theta^{\prime})f^{-1}+(c_{2}-C_{10}f^{-\rho})f^{-1},(c_{2}-C_{10}f^{-1})\theta^{\prime}+c_{2}\theta^{\prime}-C_{10}f^{-1-\rho}\bigr\}.$
We further retake $a_{0}\in\mathbb{N}_{0}$ larger again if necessary, so that for some $c_{3}>0$
$c_{2}-C_{10}f^{-\rho}\geq c_{3}\ \ \text{and}\ \ c_{2}-C_{10}f^{-1}\geq c_{3}$
on the same subset. Then we take $\mu\in(0,1]$ small enough that for some $c_{4}>0$
$c_{2}-C_{10}\theta^{\prime}>c_{4}.$
Lastly, since we have for some $c_{5}>0$
$c_{2}\theta^{\prime}-C_{10}f^{-1-\rho}\geq c_{5}\mu f^{-1-\delta}-C_{10}f^{-1-\rho},$
letting $a_{0}\in\mathbb{N}_{0}$ larger again if necessary, we are done with the proof. ∎
Proof of Proposition 2.1.
Let $\phi\in\mathcal{B}^{*}_{0}$ and $\lambda\in\mathbb{R}\setminus[m_{-},m_{+}]$ be as in the assertion, and let $\mu>0$ and $a_{0}\in\mathbb{N}_{0}$ be from Lemma 2.8. Then, taking the expectation of the inequality from Lemma 2.8 on the state $\chi_{a-2,b+2}\phi$, we obtain that for $b>a\geq a_{0}$ and $\nu\in\mathbb{N}_{0}$
$\begin{split}\bigl\|(\min\{f^{-1},\theta^{\prime}\}\Theta)^{1/2}\phi\bigr\|^{2}&\leq C_{a}\|\chi_{a-1,a+1}\phi\|^{2}+C_{\nu}2^{-\nu}\|\chi_{b-1,b+1}\phi\|^{2}\end{split}$ (2.18)
with $C_{a}$ being independent of $b,\nu$, and $C_{\nu}$ of $a,b$. Now we take the limit $b\to\infty$ in (2.18). Since the second term on the right-hand side vanishes due to the assumption $\phi\in\mathcal{B}^{*}_{0}$, it follows by Lebesgue’s monotone convergence theorem that
$\bigl\|(\bar{\chi}_{a}\min\{f^{-1},\theta^{\prime}\}\mathrm{e}^{\theta})^{1/2}\phi\bigr\|^{2}$ $\leq C_{a}\|\chi_{a-1,a+1}\phi\|^{2}$ (2.19)
with $\bar{\chi}_{a}$ being from (2.7). Next we let $\nu\to\infty$ in (2.19), and then by Lebesgue’s monotone convergence theorem again it follows that
$\bar{\chi}_{a}^{1/2}f^{-1/2}\mathrm{e}^{\mu f}\phi\in\mathcal{H}.$
This implies $\mathrm{e}^{\mu f}\phi\in L^{2}_{-1/2}\subset\mathcal{B}^{*}_{0}$, and thus we are done. ∎
2.4 A priori super-exponential decay estimate
To prove Proposition 2.2 we next present to the following key lemma.
Lemma 2.10.
Let $\lambda\gtrless m_{\pm}$, respectively, and fix any $\kappa>0$ and $\mu>0$ in the definition (2.6) of $\Theta$. Then there exist $c,C>0$ and $a_{0}\in\mathbb{N}_{0}$ such that uniformly in $b>a\geq a_{0}$ and $\nu\in\mathbb{N}_{0}$
$\pm\mathop{\mathrm{Im}}\bigl(A_{\Theta}(H-\lambda)\bigr)$ $\geq c\Theta-C\bigl(\chi_{a-1,a+1}^{2}+\chi_{b-1,b+1}^{2}\big)f^{-1}\mathrm{e}^{\theta}-(H-\lambda)\gamma(H-\lambda),$
respectively, where $\gamma=\gamma_{a,b}$ is a certain function satisfying $\mathop{\mathrm{supp}}\gamma\subset\mathop{\mathrm{supp}}\chi_{a,b}$ and $|\gamma|\leq C_{a,b}$.
Proof.
The proof is very similar to that of Lemma 2.8, and we present only key steps, omitting details of the computations. Note that here the parameters are slightly different, and we squeeze positivity in a different manner from Lemma 2.8. We only discuss the upper sign. Fix any $\lambda>m_{+}$ and $\kappa,\mu>0$ as in the assertion. We first choose $a_{0}\in\mathbb{N}_{0}$ such that for some $c_{1}>0$
$\lambda-q_{0}\geq c_{1}\ \ \text{and}\ \ \theta^{\prime}-f^{-1}\geq 2\kappa-f^{-1}\geq c_{1}$ (2.20)
uniformly in $\{x\in\mathbb{R}^{d};\ |x|\geq 2^{a_{0}}\}$, but later we may retake it larger. All below estimates are uniform in $b>a\geq a_{0}$ and $\nu\in\mathbb{N}_{0}$.
In this proof we set
$\begin{split}Q&=\bigl(f^{-1}\chi_{a,b}+|\chi_{a,b}^{\prime}|\bigr)\mathrm{e}^{\theta}+p_{j}\bigl(f^{-1}\chi_{a,b}+|\chi_{a,b}^{\prime}|\bigr)\mathrm{e}^{\theta}p_{j}\\ &\phantom{{}={}}{}+(H-\lambda)f\Theta(H-\lambda),\end{split}$ (2.21)
and now let us start to compute the left-hand side of the asserted inequality. Similarly to (2.11) and (2.12), here noting (2.20), we can deduce
$\begin{split}\mathop{\mathrm{Im}}\bigl(A_{\Theta}(H-\lambda)\bigr)&\geq\tfrac{1}{2}(\lambda-q_{0}-q_{1})(\theta^{\prime}+f^{-1})\Theta-\tfrac{1}{4}(\lambda-q_{0})(\theta^{\prime}-f^{-1})\Theta\\ &\phantom{{}={}}-\tfrac{1}{4}\mathrm{e}^{\theta/2}p_{i}h_{il}\alpha_{l}\chi_{a,b}(\lambda-q_{0})^{-1}(\theta^{\prime}-f^{-1})\alpha_{j}h_{jk}p_{k}\mathrm{e}^{\theta/2}\\ &\phantom{{}={}}-C_{1}Q,\end{split}$ (2.22)
with $h_{jk}$ defined as (2.13). To compute and bound the third term on the right-hand side of (2.22), we set
$\Xi=\chi_{a,b}(\lambda-q_{0})^{-1}(\theta^{\prime}-f^{-1}),\quad\widetilde{\Xi}=\chi_{a,b}(\lambda-\widetilde{q}_{0})^{-1}(\theta^{\prime}-f^{-1}),$
and then, similarly to (2.14), it follows that
$\begin{split}&-\tfrac{1}{4}\mathrm{e}^{\theta/2}p_{i}h_{il}\alpha_{l}\chi_{a,b}(\lambda-q_{0})^{-1}(\theta^{\prime}-f^{-1})\alpha_{j}h_{jk}p_{k}\mathrm{e}^{\theta/2}\\ &\geq-\tfrac{1}{4}\mathop{\mathrm{Re}}\bigl(\widetilde{\Xi}\mathrm{e}^{\theta/2}\delta_{ik}p_{i}p_{k}\mathrm{e}^{\theta/2}\bigr)-C_{2}Q.\end{split}$ (2.23)
We proceed similarly to (2.15), so that
$\begin{split}&-\tfrac{1}{4}\mathrm{e}^{\theta/2}p_{i}h_{il}\alpha_{l}\chi_{a,b}(\lambda-q_{0})^{-1}(\theta^{\prime}-f^{-1})\alpha_{j}h_{jk}p_{k}\mathrm{e}^{\theta/2}\\ &\geq-\tfrac{1}{4}\bigl(p_{i}\alpha_{i}+\tfrac{\mathrm{i}}{2}\alpha_{f}\theta^{\prime}\bigr)\Xi\mathrm{e}^{\theta}\bigl(\alpha_{k}p_{k}-\tfrac{\mathrm{i}}{2}\alpha_{f}\theta^{\prime}\bigr)+\tfrac{1}{4}\mathop{\mathrm{Im}}\bigl(\alpha_{i}(\partial_{i}\Xi)\mathrm{e}^{\theta}\alpha_{k}p_{k}\bigr)\\ &\phantom{{}={}}-\tfrac{1}{8}\mathop{\mathrm{Re}}\bigl(\alpha_{i}\alpha_{f}(\partial_{i}\Xi)\theta^{\prime}\mathrm{e}^{\theta}\bigr)-C_{2}Q\\ &\geq\tfrac{1}{16}\Xi\theta^{\prime 2}\mathrm{e}^{\theta}-\tfrac{1}{4}p_{i}\alpha_{i}\Xi\mathrm{e}^{\theta}\alpha_{k}p_{k}-C_{3}Q\\ &\geq c_{2}\Theta-\tfrac{1}{4}(\lambda-q_{0}-2q_{1})(\theta^{\prime}-f^{-1})\Theta-C_{4}Q.\end{split}$ (2.24)
Thus by (2.22) and (2.24) we obtain
$\begin{split}\mathop{\mathrm{Im}}\bigl(A_{\Theta}(H-\lambda)\bigr)&\geq c_{2}\Theta+(c_{1}-q_{1})f^{-1}\Theta-C_{5}Q.\end{split}$ (2.25)
We can control $Q$ as follows. In fact, similarly to (2.17), we can prove
$\begin{split}Q&\leq C_{6}f^{-1}\Theta+C_{6}\bigl(\chi_{a-1,a+1}^{2}+\chi_{b-1,b+1}^{2}\big)f^{-1}\mathrm{e}^{\theta}\\ &\phantom{{}={}}{}+C_{6}(H-\lambda)f\Theta(H-\lambda),\end{split}$ (2.26)
and hence by (2.25) and (2.26)
$\begin{split}\mathop{\mathrm{Im}}\bigl(A_{\Theta}(H-\lambda)\bigr)&\geq(c_{2}-C_{7}f^{-1})f\Theta-C_{7}\bigl(\chi_{a-1,a+1}^{2}+\chi_{b-1,b+1}^{2}\big)f^{-1}\mathrm{e}^{\theta}\\ &\phantom{{}={}}{}-C_{7}(H-\lambda)f\Theta(H-\lambda).\end{split}$
Finally by letting $a_{0}\in\mathbb{N}_{0}$ larger if necessary we are done with the proof. ∎
Proof of Proposition 2.2.
Fix $\phi\in\mathcal{B}^{*}_{0}$, $\lambda\in\mathbb{R}\setminus[m_{-},m_{+}]$ and $\kappa,\mu>0$ as in the assertion, and take $a_{0}\in\mathbb{N}_{0}$ as in Lemma 2.10. Taking the expectation of the inequality from Lemma 2.10 on the state $\chi_{a-2,b+2}\phi$, we obtain that for $b>a\geq a_{0}$ and $\nu\in\mathbb{N}_{0}$
$\begin{split}\bigl\|\Theta^{1/2}\phi\bigr\|^{2}&\leq C_{a}\|\chi_{a-1,a+1}\phi\|^{2}+C_{\nu}2^{-\nu}\|\chi_{b-1,b+1}\mathrm{e}^{\theta}\phi\|^{2}.\end{split}$ (2.27)
If we let $b\to\infty$ in (2.27), the second term on the right-hand side vanishes since $\mathrm{e}^{\kappa f}\phi\in\mathcal{B}^{*}_{0}$, and thus by Lebesgue’s monotone convergence theorem
$\bigl\|(\bar{\chi}_{a}\mathrm{e}^{\theta})^{1/2}\phi\bigr\|^{2}$ $\leq C_{a}\|\chi_{a-1,a+1}\phi\|^{2}$ (2.28)
with $\bar{\chi}_{a}$ being from (2.7). We next let $\nu\to\infty$ in (2.28), and then by Lebesgue’s monotone convergence theorem again
$\bar{\chi}_{a}^{1/2}\mathrm{e}^{(\kappa+\mu)f}\phi\in\mathcal{H}.$
This implies the assertion. ∎
2.5 Absence of super-exponentially decaying eigenstates
A key lemma for the proof of Proposition 2.3 is the following.
Lemma 2.11.
Let $\lambda\gtrless m_{\pm}$, respectively, and fix $\mu=0$ in the definition (2.6) of $\Theta$, so that $\theta=2\kappa f$, and that $\Theta$ is independent of $\nu\in\mathbb{N}_{0}$. Then there exist $c,C>0$ and $a_{0}\in\mathbb{N}_{0}$ such that uniformly in $\kappa\geq 1$ and $b>a\geq a_{0}$
$\pm\mathop{\mathrm{Im}}\bigl(A_{\Theta}(H-\lambda)\bigr)$ $\geq c\kappa^{3}\Theta-C\kappa^{2}\bigl(\chi_{a-1,a+1}^{2}+\chi_{b-1,b+1}^{2}\big)f^{-1}\mathrm{e}^{2\kappa f}$
$\phantom{={}}{}-(H-\lambda)\gamma(H-\lambda),$
respectively, where $\gamma=\gamma_{a,b,\kappa}$ is a certain function satisfying $\mathop{\mathrm{supp}}\gamma\subset\mathop{\mathrm{supp}}\chi_{a,b}$ and $|\gamma|\leq C_{a,b,\kappa}$.
Proof.
Again the proof is very similar to those of Lemmas 2.8 and 2.10, and we present only key steps. Only the upper sign is discussed. Fix any $\lambda>m_{+}$, and let $\mu=0$. We first choose $a_{0}\in\mathbb{N}_{0}$ such that for some $c_{1}>0$
$\lambda-q_{0}\geq c_{1}\ \ \text{and}\ \ \theta^{\prime}-f^{-1}=2\kappa-f^{-1}\geq c_{1}\kappa$ (2.29)
uniformly in $\{x\in\mathbb{R}^{d};\ |x|\geq 2^{a_{0}}\}$ and $\kappa\geq 1$, but later we may retake it larger. All below estimates are uniform in $\kappa\geq 1$ and $b>a\geq a_{0}$.
In this proof we set
$\begin{split}Q&=\kappa^{2}\bigl(f^{-1}\chi_{a,b}+|\chi_{a,b}^{\prime}|\bigr)\mathrm{e}^{2\kappa f}+p_{j}\bigl(f^{-1}\chi_{a,b}+|\chi_{a,b}^{\prime}|\bigr)\mathrm{e}^{2\kappa f}p_{j}\\ &\phantom{{}={}}{}+\kappa(H-\lambda)f\Theta(H-\lambda),\end{split}$ (2.30)
and then, similarly to (2.11) and (2.12), or (2.22), noting (2.29), we can deduce
$\begin{split}\mathop{\mathrm{Im}}\bigl(A_{\Theta}(H-\lambda)\bigr)&\geq\tfrac{1}{2}(\lambda-q_{0}-q_{1})(2\kappa+f^{-1})\Theta-\tfrac{1}{4}(\lambda-q_{0})(2\kappa-f^{-1})\Theta\\ &\phantom{{}={}}-\tfrac{1}{4}\mathrm{e}^{\kappa f}p_{i}h_{il}\alpha_{l}\chi_{a,b}(\lambda-q_{0})^{-1}(2\kappa-f^{-1})\alpha_{j}h_{jk}p_{k}\mathrm{e}^{\kappa f}\\ &\phantom{{}={}}-C_{1}Q,\end{split}$ (2.31)
with $h_{jk}$ being from (2.13). Set
$\Xi=\chi_{a,b}(\lambda-q_{0})^{-1}(2\kappa-f^{-1}),\quad\widetilde{\Xi}=\chi_{a,b}(\lambda-\widetilde{q}_{0})^{-1}(2\kappa-f^{-1}),$
and then, similarly to (2.14) or (2.23), the third term on the right-hand side of (2.31) is bounded as
$\begin{split}&-\tfrac{1}{4}\mathrm{e}^{\kappa f}p_{i}h_{il}\alpha_{l}\chi_{a,b}(\lambda-q_{0})^{-1}(2\kappa-f^{-1})\alpha_{j}h_{jk}p_{k}\mathrm{e}^{\kappa f}\\ &\geq-\tfrac{1}{4}\mathop{\mathrm{Re}}\bigl(\widetilde{\Xi}\mathrm{e}^{\kappa f}\delta_{ik}p_{i}p_{k}\mathrm{e}^{\kappa f}\bigr)-C_{2}Q.\end{split}$
We proceed similarly to (2.15) or (2.24) as
$\begin{split}&-\tfrac{1}{4}\mathrm{e}^{\kappa f}p_{i}h_{il}\alpha_{l}\chi_{a,b}(\lambda-q_{0})^{-1}(2\kappa-f^{-1})\alpha_{j}h_{jk}p_{k}\mathrm{e}^{\kappa f}\\ &\geq-\tfrac{1}{4}\bigl(p_{i}\alpha_{i}+\mathrm{i}\kappa\alpha_{f}\bigr)\Xi\mathrm{e}^{2\kappa f}\bigl(\alpha_{k}p_{k}-\mathrm{i}\kappa\alpha_{f}\bigr)+\tfrac{1}{4}\mathop{\mathrm{Im}}\bigl(\alpha_{i}(\partial_{i}\Xi)\mathrm{e}^{2\kappa f}\alpha_{k}p_{k}\bigr)\\ &\phantom{{}={}}-\tfrac{1}{4}\kappa\mathop{\mathrm{Re}}\bigl(\alpha_{i}\alpha_{f}(\partial_{i}\Xi)\mathrm{e}^{2\kappa f}\bigr)-C_{2}Q\\ &\geq\tfrac{1}{4}\kappa^{2}\Xi\mathrm{e}^{2\kappa f}-\tfrac{1}{4}p_{i}\alpha_{i}\Xi\mathrm{e}^{2\kappa f}\alpha_{k}p_{k}-C_{3}Q\\ &\geq c_{2}\kappa^{3}\Theta-\tfrac{1}{4}(\lambda-q_{0}-2q_{1})(2\kappa-f^{-1})\Theta-C_{4}Q.\end{split}$ (2.32)
Thus by (2.31) and (2.32)
$\begin{split}\mathop{\mathrm{Im}}\bigl(A_{\Theta}(H-\lambda)\bigr)&\geq c_{2}\kappa^{3}\Theta+(c_{1}-q_{1})f^{-1}\Theta-C_{5}Q.\end{split}$ (2.33)
Similarly to (2.17) or (2.26), we have
$\begin{split}Q&\leq C_{6}\kappa^{2}f^{-1}\Theta+C_{6}\kappa^{2}\bigl(\chi_{a-1,a+1}^{2}+\chi_{b-1,b+1}^{2}\big)f^{-1}\mathrm{e}^{\theta}\\ &\phantom{{}={}}{}+C_{6}\kappa(H-\lambda)f\Theta(H-\lambda),\end{split}$ (2.34)
so that by (2.33) and (2.34)
$\begin{split}\mathop{\mathrm{Im}}\bigl(A_{\Theta}(H-\lambda)\bigr)&\geq\kappa^{3}(c_{2}-C_{7}\kappa^{-1}f^{-1})f\Theta-C_{7}\kappa^{2}\bigl(\chi_{a-1,a+1}^{2}+\chi_{b-1,b+1}^{2}\big)f^{-1}\mathrm{e}^{\theta}\\ &\phantom{{}={}}{}-C_{7}\kappa(H-\lambda)f\Theta(H-\lambda).\end{split}$
Letting $a_{0}\in\mathbb{N}_{0}$, we are done with the proof. ∎
Proof of Proposition 2.3.
Let $\phi\in\mathcal{B}^{*}_{0}$ and $\lambda\in\mathbb{R}\setminus[m_{-},m_{+}]$ be as in the assertion. Let $a_{0}\in\mathbb{N}_{0}$ be from Lemma 2.11, and take the expectation of the inequality from Lemma 2.11 on the state $\chi_{a-2,b+2}\phi$. Then for any $\kappa\geq 1$ and $b>a\geq a_{0}$
$\kappa\|\chi_{a,b}^{1/2}\mathrm{e}^{\kappa f}\phi\|^{2}\leq C\|\chi_{a-1,a+1}\mathrm{e}^{\kappa f}\phi\|^{2}+C2^{-b}\|\chi_{b-1,b+1}\mathrm{e}^{\kappa f}\phi\|^{2}.$ (2.35)
By the assumption the second term to the right of (2.35) vanishes as $b\to\infty$, and hence by Lebesgue’s monotone convergence theorem
$\kappa\|\bar{\chi}_{a}^{1/2}\mathrm{e}^{\kappa f}\|^{2}$ $\leq C\|\chi_{b-1,b+1}\mathrm{e}^{\kappa f}\phi\|^{2},$
or
$\kappa\|\bar{\chi}_{a}^{1/2}\mathrm{e}^{\kappa(f-2^{b+2})}\phi\|^{2}$ $\leq C\|\chi_{b-1,b+1}\phi\|^{2}.$ (2.36)
Now we assume $\bar{\chi}_{b+2}\phi\neq 0$ as a function on $\mathbb{R}^{d}$. Then the left-hand side of (2.36) grows exponentially as $\kappa\to\infty$ whereas the right-hand side remains bounded. This is a contradiction, and thus $\bar{\chi}_{b+2}\phi=0$. Then by the unique continuation property [10] we conclude that $\phi=0$ as a function on $\mathbb{R}^{d}$. We are done. ∎
3 Proof of LAP bounds
In this section prove Theorem 1.5. The proof again relies on commutator arguments [16] with conjugate operator $A_{\Theta}$ from (2.1), however, with a different weight function $\Theta$ from that for Rellich’s theorem. We in fact use
$\Theta=\Theta_{a,\nu}=\bar{\chi}_{a}\theta,\quad\theta=\theta_{\nu}=\int_{0}^{f/2^{\nu}}(1+\tau)^{-1-\delta}\,\mathrm{d}\tau=\delta^{-1}\bigl(1-(1+f/2^{\nu})^{-\delta}\bigr),$ (3.1)
depending on parameters $a,\nu\in\mathbb{N}_{0}$, see (2.7) for the definition of $\bar{\chi}_{a}$. Here $\delta\in(0,\rho)$ is an arbitrarily fixed constant, and the dependence on it is suppressed since we shall not retake it in this section.
In Section 3.1 we present preliminary commutator computations. In Section 3.2 we prove Theorem 1.5 by contradiction to Rellich’s theorem. Throughout the section we assume Assumption 1.1.
3.1 Preliminaries
3.1.1 Properties of weight function
Here we present a key commutator estimate for the proof of Theorem 1.5, but before that let us recall some elementary properties of the function $\theta$ from (3.1). As in the previous section we denote the derivatives in $f$ by primes, such as
$\theta^{\prime}=2^{-\nu}(1+f/2^{\nu})^{-1-\delta},\quad\theta^{\prime\prime}=-2^{-2\nu}(1+\delta)(1+f/2^{\nu}{})^{-2-\delta}.$ (3.2)
Lemma 3.1.
There exist $c,C>0$ such that uniformly in $\nu\in\mathbb{N}_{0}$
$c/2^{\nu}{}\leq\theta\leq\min\{C,f/2^{\nu}\},\quad c(\min\{2^{\nu},f\})^{\delta}f^{-1-\delta}\theta\leq\theta^{\prime}\leq f^{-1}\theta.$
In addition, for any $k=2,3,\dots$ there exists $C_{k}$ such that uniformly in $\nu\in\mathbb{N}_{0}$
$0\leq(-1)^{k-1}\theta^{(k)}\leq C_{k}f^{-k}\theta.$
Proof.
See [16] for the proof. ∎
The weight $\theta^{\prime}$ conforms with the $\mathcal{B}^{*}$-norm in the following manner.
Lemma 3.2.
There exist $c,C>0$ such that for any $\psi\in\mathcal{B}^{*}$
$c\|\psi\|_{\mathcal{B}^{*}}\leq\sup_{\nu\in\mathbb{N}_{0}}\|\theta^{\prime 1/2}\psi\|\leq C\|\psi\|_{\mathcal{B}^{*}}.$
Proof.
The former inequality is obvious since there exists $C_{1}>0$ such that uniformly in $\nu\in\mathbb{N}_{0}$
$2^{-\nu}F_{\nu}\leq C_{1}\theta^{\prime},$
where $F_{\nu}$ is from (1.9). On the other hand, as for the latter, by noting the expression (3.2) we can decompose and bound the norm for any $\nu\in\mathbb{N}_{0}$ as
$\|\theta^{\prime 1/2}\psi\|$ $\leq\sum_{\mu=0}^{\nu-1}\|F_{\mu}\theta^{\prime 1/2}\psi\|+\sum_{\mu=\nu}^{\infty}\|F_{\mu}\theta^{\prime 1/2}\psi\|$
$\leq\sum_{\mu=0}^{\nu-1}2^{-\nu/2}\|F_{\mu}\psi\|+\sum_{\mu=\nu}^{\infty}2^{\delta\nu/2}\|F_{\mu}f^{-(1+\delta)/2}\psi\|$
$\leq\sum_{\mu=0}^{\nu-1}2^{-(\nu-\mu)/2}2^{-\mu/2}\|F_{\mu}\psi\|+\sum_{\mu=\nu}^{\infty}2^{-\delta(\mu-\nu)/2}2^{-\mu/2}\|F_{\mu}\psi\|$
$\leq\|\psi\|_{\mathcal{B}^{*}}\Biggl(\sum_{\kappa=1}^{\nu}2^{-\kappa/2}+\sum_{\kappa=0}^{\infty}2^{-\delta\kappa/2}\Biggr).$
This in fact implies the latter bound. ∎
3.1.2 Commutator estimate
Now we state and prove the main commutator inequality of the section. Recall notation $\sigma$ defined right after (1.12).
Lemma 3.3.
Let $I\subset\mathbb{R}\setminus[m_{-},m_{+}]$ be a compact interval. Then there exist $c,C>0$ and $a,b\in\mathbb{N}_{0}$ such that uniformly in $z\in I_{\pm}$ and $\nu\in\mathbb{N}_{0}$
$\sigma\mathop{\mathrm{Im}}\bigl(A_{\Theta}(H-z)\bigr)\geq c\theta^{\prime}-C\chi_{b}^{2}\theta-\mathop{\mathrm{Re}}(\gamma(H-z))-C(H-z)^{*}(H-z),$
respectively, where $\gamma=\gamma_{z,\nu}$ is a uniformly bounded $n\times n$ matrix-valued function, i.e., $|\gamma|\leq C$ independently of $z\in I_{\pm}$ and $\nu\in\mathbb{N}_{0}$.
Proof.
We discuss only the case with $\sigma=+$ and the upper sign. Fix a compact interval $I\subset(m_{+},\infty)$, and choose $a\in\mathbb{N}_{0}$ such that for some $c_{1}>0$
$\lambda-q_{0}\geq c_{1}\ \ \text{uniformly in }\lambda\in I\text{ and }|x|\geq 2^{a}.$
The below estimates are all uniform in $z=\lambda+\mathrm{i}\Gamma\in I_{+}$ and $\nu\in\mathbb{N}_{0}$. In this proof for notational simplicity we let
$Q=f^{-1-\rho}\theta+p_{j}f^{-1-\rho}\theta p_{j}+(H-z)^{*}(H-z).$
Now we compute and bound the left-hand side of the asserted inequality. By Lemmas 2.4 and 3.1 and the Cauchy–Schwarz inequality we first have
$\begin{split}\mathop{\mathrm{Im}}\bigl(A_{\Theta}(H-z)\bigr)&=\mathop{\mathrm{Re}}\bigl(\alpha_{f}\Theta^{\prime}p_{f}\bigr)+\mathop{\mathrm{Re}}\bigl(\alpha_{j}f^{-1}\Theta\ell_{jk}p_{k}\bigr)-\bigl(\partial_{f}(q_{0}+q_{1})\bigr)\Theta\\ &\phantom{{}={}}{}+(\Delta f)q_{2}\Theta+q_{2}\Theta^{\prime}-2\mathop{\mathrm{Im}}(q_{2}\Theta p_{f})-\Gamma A_{\Theta}\\ &\geq\mathop{\mathrm{Re}}\bigl(\alpha_{f}\bar{\chi}_{a}\theta^{\prime}p_{f}\bigr)+\mathop{\mathrm{Re}}\bigl(\alpha_{j}f^{-1}\Theta\ell_{jk}p_{k}\bigr)-\Gamma A_{\Theta}-C_{1}Q.\end{split}$ (3.3)
The first and second terms on the right-hand side of (3.3) combine, similarly to (2.12), as
$\begin{split}&\mathop{\mathrm{Re}}\bigl(\alpha_{f}\bar{\chi}_{a}\theta^{\prime}p_{f}\bigr)+\mathop{\mathrm{Re}}\bigl(\alpha_{j}f^{-1}\Theta\ell_{jk}p_{k}\bigr)\\ &=\tfrac{1}{2}(\lambda-q)\bar{\chi}_{a}(\theta^{\prime}+f^{-1}\theta)+\tfrac{1}{2}\mathop{\mathrm{Re}}\bigl(\bar{\chi}_{a}(\theta^{\prime}+f^{-1}\theta)(H-z)\bigr)\\ &\phantom{={}}{}{}-\tfrac{1}{2}\mathop{\mathrm{Re}}\bigl(\bar{\chi}_{a}(f^{-1}\theta-\theta^{\prime})\alpha_{j}h_{jk}p_{k}\bigr)\\ &\geq\tfrac{1}{2}(\lambda-q_{0}-q_{1})\bar{\chi}_{a}(\theta^{\prime}+f^{-1}\theta)-\tfrac{1}{4}(\lambda-q_{0})\bar{\chi}_{a}(f^{-1}\theta-\theta^{\prime})\\ &\phantom{={}}{}-\tfrac{1}{4}p_{i}h_{il}\alpha_{l}(\lambda-q_{0})^{-1}\bar{\chi}_{a}(f^{-1}\theta-\theta^{\prime})\alpha_{j}h_{jk}p_{k}-C_{2}Q,\end{split}$ (3.4)
with $h_{jk}$ being from (2.13). To bound the third term on the right-hand side of (3.4), we set
$\Xi=(\lambda-q_{0})^{-1}\bar{\chi}_{a}(f^{-1}\theta-\theta^{\prime}),\quad\widetilde{\Xi}=(\lambda-\widetilde{q}_{0})^{-1}\bar{\chi}_{a}(f^{-1}\theta-\theta^{\prime})$
for short, and proceed similarly to (2.14) and (2.15) as
$\begin{split}&-\tfrac{1}{4}p_{i}h_{il}\alpha_{l}(\lambda-q_{0})^{-1}\bar{\chi}_{a}(f^{-1}\theta-\theta^{\prime})\alpha_{j}h_{jk}p_{k}\\ &=-\tfrac{1}{4}\mathop{\mathrm{Re}}\bigl(\widetilde{\Xi}h_{ij}h_{jk}p_{i}p_{k}\bigr)-\tfrac{1}{4}\mathop{\mathrm{Im}}\bigl(\alpha_{l}(\partial_{i}h_{il}\Xi h_{jk})\alpha_{j}p_{k}\bigr)\\ &\geq-\tfrac{1}{4}\mathop{\mathrm{Re}}\bigl(\widetilde{\Xi}\delta_{ik}p_{i}p_{k}\bigr)-C_{3}Q\\ &=-\tfrac{1}{4}p_{i}\alpha_{i}\Xi\alpha_{k}p_{k}+\tfrac{1}{4}\mathop{\mathrm{Im}}\bigl(\alpha_{i}(\partial_{i}\Xi)\alpha_{k}p_{k}\bigr)-C_{3}Q\\ &\geq-\tfrac{1}{4}(\lambda-q_{0}-2q_{1})\bar{\chi}_{a}(f^{-1}\theta-\theta^{\prime})-C_{4}Q.\end{split}$ (3.5)
Hence by (3.4) and (3.5)
$\begin{split}\mathop{\mathrm{Re}}\bigl(\alpha_{f}\bar{\chi}_{a}\theta^{\prime}p_{f}\bigr)+\mathop{\mathrm{Re}}\bigl(\alpha_{j}f^{-1}\Theta\ell_{jk}p_{k}\bigr)&\geq(\lambda-q_{0}-q_{1})\bar{\chi}_{a}\theta^{\prime}-C_{5}Q\\ &\geq(c_{1}-q_{1})\theta^{\prime}-C_{6}Q.\end{split}$ (3.6)
As for the third term on the right-hand side of (3.3), we use the Cauchy–Schwarz inequality and Lemma 2.6 with $\Xi=1$ to deduce
$\begin{split}-\Gamma A_{\Theta}&\geq-C_{7}\Gamma p_{j}p_{j}-C_{7}\Gamma\\ &\geq-C_{7}\Gamma\bigl(p_{j}p_{j}-(z-q)^{*}(z-q)\bigr)-C_{8}\Gamma\\ &=-2C_{7}\Gamma\mathop{\mathrm{Re}}\bigl((z-q)^{*}(H-z)\bigr)-C_{7}\Gamma(H-z)^{*}(H-z)\\ &\phantom{={}}{}+C_{8}\mathop{\mathrm{Im}}(H-z)\\ &\geq-\mathop{\mathrm{Re}}(\gamma(H-z))-C_{9}Q,\end{split}$ (3.7)
where we have set
$\gamma=2C_{7}\Gamma(z-q)^{*}+\mathrm{i}C_{8}$
Therefore by (3.3), (3.6) and (3.7) it follows that
$\mathop{\mathrm{Im}}\bigl(A_{\Theta}(H-z)\bigr)\geq(c_{1}-q_{1})\theta^{\prime}-\mathop{\mathrm{Re}}(\gamma(H-z))-C_{10}Q,$ (3.8)
Finally we can show by Lemmas 2.6 and 3.1 that
$\begin{split}Q&\leq C_{11}f^{-1-\rho}\theta+C_{11}(H-z)^{*}(H-z)\\ &\leq C_{12}f^{-(\rho-\delta)}\theta^{\prime}+C_{11}(H-z)^{*}(H-z),\end{split}$ (3.9)
which, combined with (3.8), implies
$\mathop{\mathrm{Im}}\bigl(A_{\Theta}(H-z)\bigr)$ $\geq\bigl(c_{1}-q_{1}-C_{13}f^{-(\rho-\delta)}\bigr)\theta^{\prime}-\mathop{\mathrm{Re}}(\gamma(H-z))$
$\phantom{={}}{}+C_{13}(H-z)^{*}(H-z).$
Thus by letting $b\in\mathbb{N}_{0}$ be large enough we obtain the assertion. ∎
3.2 Contradiction to Rellich’s theorem
Here we provide the proof of the LAP bounds.
Proof of Theorem 1.5.
Let $I\subset\mathbb{R}\setminus[m_{-},m_{+}]$ be a compact interval, and fix $a,b\in\mathbb{N}_{0}$ as the assertion of Lemma 3.3,
Step I. We first prove that there exists $C_{1}>0$ such that for any $\phi=R(z)\psi$ with $z\in I_{\pm}$ and $\psi\in\mathcal{B}$
$\|p_{1}\phi\|_{\mathcal{B}^{*}}+\dots+\|p_{d}\phi\|_{\mathcal{B}^{*}}\leq C_{1}\bigl(\|\phi\|_{\mathcal{B}^{*}}+\|\psi\|_{\mathcal{B}^{*}}\bigr).$ (3.10)
In fact, by Lemma 2.6 with $\Xi=\theta^{\prime}$ we have
$p_{j}\theta^{\prime}p_{j}$ $=(z-q)^{*}(z-q)\theta^{\prime}-\mathop{\mathrm{Im}}\bigl(\alpha_{j}(\partial_{j}\theta^{\prime})(z-q)\bigr)+\tfrac{1}{2}(\Delta\theta^{\prime})$
$\phantom{={}}{}+\mathop{\mathrm{Re}}\bigl(\bigl(2(z-q)^{*}\theta^{\prime}+\mathrm{i}\alpha_{j}(\partial_{j}\theta^{\prime})\bigr)(H-z)\bigr)+(H-z)^{*}\theta^{\prime}(H-z),$
and, if we take the expectation of this inequality on the state $\phi=R(z)\psi$ and use the Cauchy–Schwarz inequality, it follows that
$\|\theta^{\prime 1/2}p_{1}\phi\|^{2}+\dots+\|\theta^{\prime 1/2}p_{d}\phi\|^{2}$ $\leq C_{2}\bigl(\|\theta^{\prime 1/2}\phi\|^{2}+\|\theta^{\prime 1/2}\psi\|^{2}\bigr).$
Hence by Lemma 3.2 we obtain the claim (3.10).
Step II. Next we claim that there exist $C_{3},C_{4}>0$ such that uniformly in $\nu\in\mathbb{N}_{0}$ and $\phi=R(z)\psi$ with $z\in I_{\pm}$ and $\psi\in\mathcal{B}$
$\|\theta^{\prime}{}^{1/2}\phi\|^{2}$ $\leq C_{3}\bigl(\|\psi\|_{\mathcal{B}}^{2}+\|\chi_{b}\theta^{1/2}\phi\|^{2}\bigr),$ (3.11)
$\|\phi\|_{\mathcal{B}^{*}}^{2}$ $\leq C_{4}\bigl(\|\psi\|_{\mathcal{B}}^{2}+\|\chi_{b}\phi\|^{2}\bigr).$ (3.12)
In fact, take the expectation of the inequality from Lemma 3.3 on the state $\phi=R(z)\psi$, and use the $\mathcal{B}$-$\mathcal{B}^{*}$ duality, the Cauchy–Schwarz inequality and (3.10), and then we can immediately deduce (3.11). In addition, (3.11) directly follows from (3.11) and Lemmas 3.1 and 3.2.
Step III. Finally we prove the assertion. By (3.10) it suffices to show that uniformly in $\phi=R(z)\psi$ with $z\in I_{\pm}$ and $\psi\in\mathcal{B}$
$\|\phi\|_{\mathcal{B}^{*}}\leq C_{5}\|\psi\|_{\mathcal{B}}.$ (3.13)
We assume that this is false, and choose $z_{k}\in I_{\pm}$ and $\psi_{k}\in\mathcal{B}$ such that
$\lim_{k\to\infty}\|\psi_{k}\|_{\mathcal{B}}=0,\quad\|\phi_{k}\|_{\mathcal{B}^{*}}=1;\quad\phi_{k}=R(z_{k})\psi_{k}.$ (3.14)
We first note that we may let $(z_{k})_{k\in\mathbb{N}}$ be convergent to a point in $I\subset\mathbb{R}$, i.e., for some $\lambda\in I$
$\lambda=\lim_{k\to\infty}z_{k}\in I.$ (3.15)
In fact, since
$\|\phi_{k}\|_{\mathcal{B}^{*}}\leq\|\phi_{k}\|\leq\|R(z_{k})\|_{\mathcal{L}(\mathcal{H})}\|\psi_{k}\|\leq C_{6}|\mathop{\mathrm{Im}}z_{k}|^{-1}\|\psi_{k}\|_{\mathcal{B}},$
we have $\mathop{\mathrm{Im}}z_{k}\to 0$ as $k\to\infty$ due to (3.14). Then by compactness of $I$, we can choose a convergent subsequence of $(\mathop{\mathrm{Re}}z_{k})_{k\in\mathbb{N}}$.
Next, we show that for any $s>1/2$ by choosing a further subsequence we may let $(\phi_{k})_{k\in\mathbb{N}}$ be convergent in $L^{2}_{-s}$, i.e., for some $\phi\in L^{2}_{-s}$
$\phi=\lim_{k\to\infty}\phi_{k}\in L^{2}_{-s}.$ (3.16)
In fact, choose any $s^{\prime}\in(1/2,s)$, and then by (3.14) and (3.10) the sequence $(f^{-s^{\prime}}\phi_{k})_{k\in\mathbb{N}}$ is uniformly bounded in $H^{1}(\mathbb{R}^{d};\mathbb{C}^{n})$. Since the multiplication operator
$f^{-(s-s^{\prime})}\colon H^{1}(\mathbb{R}^{d};\mathbb{C}^{n})\to L^{2}(\mathbb{R}^{d};\mathbb{C}^{n})$
is compact, we obtain the claim (3.16).
By (3.14), (3.15) and (3.16) it follows that
$(H-\lambda)\phi=0\text{ in the distributional sense}.$ (3.17)
Furthermore, we can verify $\phi\in\mathcal{B}^{*}_{0}$, so that $\phi=0$ by Theorem 1.3. In fact, by (3.11) we have
$\|\theta^{\prime}{}^{1/2}\phi_{k}\|^{2}\leq C_{3}\bigl(\|\psi_{k}\|_{\mathcal{B}}^{2}+\|\chi_{b}\theta^{1/2}\phi_{k}\|^{2}\bigr).$
If we let $s=(1+\delta)/2$ and take the limit $k\to\infty$, we obtain by (3.16), (3.14) and Lemma 3.1 that for any $\nu\in\mathbb{N}_{0}$
$\begin{split}\|\theta^{\prime}{}^{1/2}\phi\|&\leq C_{7}\|\chi_{b}\theta^{1/2}\phi\|\leq C_{7}2^{-\nu/2}\|\chi_{b}f^{1/2}\phi\|.\end{split}$ (3.18)
Letting $\nu\to\infty$ in (3.18), we obtain $\phi\in\mathcal{B}^{*}_{0}$, and then conclude $\phi=0$.
However, this is a contradiction, because by (3.12) we have
$1=\|\phi_{k}\|_{\mathcal{B}^{*}}^{2}\leq C_{4}\bigl(\|\psi_{k}\|_{\mathcal{B}}^{2}+\|\chi_{b}\phi_{k}\|^{2}\big),$
and, as $k\to\infty$, the right-hand side converges to 0 by (3.14) and (3.16) and $\phi=0$. Therefore (3.13) holds true. ∎
4 Proof of radiation condition bounds, and applications
4.1 Commutator method for radiation condition bounds
Here we prove Theorem 1.9. The proof, again, depends on commutator arguments, however, it is fairly different from Ito–Skibsted [16]. Theorem 1.9 is almost trivial from the following bounds.
Proposition 4.1.
Suppose Assumption 1.7, let $I\subset\mathbb{R}\setminus[m_{-},m_{+}]$ be a compact interval, and let $\kappa\in(0,1/2)$. Then there exists $C>0$ such that for any $z\in I_{\pm}$
$\pi_{\mp}^{2}f^{-1+2\kappa}+(p_{f}-\alpha_{f}(z-q_{0}))^{*}f^{-1+2\kappa}(p_{f}-\alpha_{f}(z-q_{0}))$
$\leq Cf^{-1-\rho+2\kappa}+C(H-z)^{*}f^{1+2\kappa}(H-z),$
respectively.
We split the proof of Proposition 4.1 into two steps, and for the first step we use a commutator method. The weight function of the section is given by
$\Theta=\Theta^{\kappa}_{a}=\bar{\chi}_{a}f^{2\kappa}$
with parameters $a\in\mathbb{N}$ and $\kappa>0$. Recall notation $\sigma$ defined right after (1.12).
Lemma 4.2.
Suppose Assumption 1.7, let $I\subset\mathbb{R}\setminus[m_{-},m_{+}]$ be a compact interval, and let $\kappa>0$. There exist $c,C>0$ and $a\in\mathbb{N}$ such that for any $z\in I_{\pm}$
$\mathop{\mathrm{Im}}\bigl((\sigma p_{f}\mp(z-q_{0}))^{*}\Theta(H-z)\bigr)$
$\geq c\pi_{\mp}^{2}f^{-1+2\kappa}-\sigma\bigl(\tfrac{1}{2}-\kappa\bigr)\mathop{\mathrm{Re}}\bigl(f^{-1}\Theta(\alpha_{f}p_{f}-z+q_{0})\bigr)$
$\phantom{{}={}}{}-Cf^{-1-\rho+2\kappa}-C(H-z)^{*}f^{1+2\kappa}(H-z),$
respectively. In particular, it follows that for any $\epsilon>0$ there exist $c^{\prime},C^{\prime}>0$ such that for any $z\in I_{\pm}$
$c^{\prime}\pi_{\mp}^{2}f^{-1+2\kappa}-\sigma\bigl(\tfrac{1}{2}-\kappa\bigr)\mathop{\mathrm{Re}}\bigl(f^{-1}\Theta(\alpha_{f}p_{f}-z+q_{0})\bigr)$
$\leq C^{\prime}f^{-1-\rho+2\kappa}+C^{\prime}(H-z)^{*}f^{1+2\kappa}(H-z)$
$\phantom{{}={}}{}-\epsilon(p_{f}-\alpha_{f}(z-q_{0}))^{*}f^{-1+2\kappa}(p_{f}-\alpha_{f}(z-q_{0})),$
respectively.
Proof.
The latter inequality follows from the former, since by Assumption 1.7 and the Cauchy–Schwarz inequality
$\mathop{\mathrm{Im}}\bigl((\sigma p_{f}\mp(z-q_{0}))^{*}\Theta(H-z)\bigr)$
$=\sigma\mathop{\mathrm{Im}}\bigl((p_{f}-\alpha_{f}(z-q_{0}))^{*}\Theta(H-z)\bigr)\mp 2\mathop{\mathrm{Im}}\bigl(\pi_{\mp}(z^{*}-q_{0})\Theta(H-z)\bigr)$
$\phantom{{}={}}{}+\sigma\mathop{\mathrm{Im}}\bigl(\alpha_{f}(q_{0}-\widetilde{q}_{0})\Theta(H-z)\bigr)$
$\leq\epsilon(p_{f}-\alpha_{f}(z-q_{0}))^{*}f^{-1+2\kappa}(p_{f}-\alpha_{f}(z-q_{0}))+\tfrac{1}{2}c^{\prime}\pi_{\mp}^{2}f^{-1+2\kappa}$
$\phantom{{}={}}{}+C_{1}f^{-1-\rho+2\kappa}+C_{1}(H-z)^{*}f^{1+2\kappa}(H-z).$
Hence it suffices to show the former inequality of the assertion. Fix any compact interval $I\subset\mathbb{R}\setminus[m_{-},m_{+}]$ and $\kappa>0$. Choose $a\in\mathbb{N}$ such that
$\sigma(\lambda-q_{0})\geq c_{1}>0\ \ \text{uniformly in }\lambda\in I\text{ and }|x|\geq 2^{a}.$
The below estimates are uniform in $z=\lambda+\mathrm{i}\Gamma\in I_{\pm}$, and we set for short
$Q=f^{-1-\rho+2\kappa}+p_{j}f^{-1-\rho+2\kappa}p_{j}+(H-z)^{*}f^{1+2\kappa}(H-z).$
Note that by Lemma 2.6 and the Cauchy–Schwarz inequality we can see that
$Q\geq-C_{2}f^{-1-\rho+2\kappa}-C_{2}(H-z)^{*}f^{1+2\kappa}(H-z),$
and thus $Q$ is in fact negligible.
Now we start to compute the left-hand side of the former assertion. We expand it, and use (2.3), (2.13), Assumption 1.7, the Cauchy–Schwarz inequality and Lemma 2.4 to have
$\begin{split}&\mathop{\mathrm{Im}}\bigl((\sigma p_{f}\mp(z-q_{0}))^{*}\Theta(H-z)\bigr)\\ &=\sigma\mathop{\mathrm{Im}}\bigl(p_{f}\Theta(H-\lambda)\bigr)\mp\tfrac{1}{2}\mathop{\mathrm{Im}}\bigl((2\lambda-q_{0}-\widetilde{q}_{0})\Theta H\bigr)-\sigma\Gamma\mathop{\mathrm{Re}}(p_{f}^{*}\Theta)\\ &\phantom{{}={}}{}\pm\tfrac{1}{2}\Gamma(2\lambda-q_{0}-\widetilde{q}_{0})\Theta\pm\tfrac{1}{2}\mathop{\mathrm{Im}}\bigl((q_{0}-\widetilde{q}_{0})\Theta(H-z)\bigr)\pm\Gamma\mathop{\mathrm{Re}}(\Theta(H-z))\\ &\geq\tfrac{1}{2}\sigma\mathop{\mathrm{Im}}(A_{\Theta}H)-\tfrac{1}{2}\sigma\mathop{\mathrm{Re}}\bigl(((\Delta f)\Theta+\Theta^{\prime})(H-z)\bigr)\\ &\phantom{{}={}}{}\mp\tfrac{1}{4}\sigma(2\lambda\alpha_{f}-\alpha_{f}q_{0}-q_{0}\alpha_{f})\Theta^{\prime}-\sigma\Gamma\mathop{\mathrm{Re}}\bigl(\alpha_{f}\Theta\alpha_{j}p_{j}\bigr)\\ &\phantom{{}={}}{}+\sigma\Gamma\mathop{\mathrm{Re}}\bigl(\alpha_{f}\Theta\alpha_{j}\ell_{jk}p_{k}\bigr)\pm\tfrac{1}{2}\Gamma(2\lambda-q_{0}-\widetilde{q}_{0})\Theta\mp C_{3}\sigma\Gamma f^{-\rho}\Theta-C_{3}Q\\ &\geq\tfrac{1}{2}\sigma\mathop{\mathrm{Re}}(\alpha_{f}\Theta^{\prime}p_{f})+\tfrac{1}{2}\sigma\mathop{\mathrm{Re}}(\alpha_{j}f^{-1}\Theta\ell_{jk}p_{k}\bigr)\mp\tfrac{1}{4}\sigma(2\lambda\alpha_{f}-\alpha_{f}q_{0}-q_{0}\alpha_{f})\Theta^{\prime}\\ &\phantom{{}={}}{}-\tfrac{1}{2}\sigma\Gamma\Theta(2\lambda\alpha_{f}-\alpha_{f}q_{0}-q_{0}\alpha_{f})+\sigma\Gamma\mathop{\mathrm{Re}}\bigl(\alpha_{l}\alpha_{j}\Theta(\partial_{l}f)\ell_{jk}p_{k}\bigr)\\ &\phantom{{}={}}{}\pm\tfrac{1}{2}\Gamma(2\lambda-q_{0}-\widetilde{q}_{0})\Theta\mp C_{4}\sigma\Gamma f^{-\rho}\Theta-C_{4}Q.\end{split}$ (4.1)
We combine the first to the third terms of (4.1) by using (2.3), Assumption 1.7 and the Cauchy–Schwarz inequality as
$\begin{split}&\tfrac{1}{2}\sigma\mathop{\mathrm{Re}}(\alpha_{f}\Theta^{\prime}p_{f})+\tfrac{1}{2}\sigma\mathop{\mathrm{Re}}(\alpha_{j}f^{-1}\Theta\ell_{jk}p_{k}\bigr)\mp\tfrac{1}{4}\sigma(2\lambda\alpha_{f}-\alpha_{f}q_{0}-q_{0}\alpha_{f})\Theta^{\prime}\\ &\geq 2\sigma\kappa\pi_{\mp}(\lambda-q_{0})\pi_{\mp}f^{-1}\Theta-\sigma\bigl(\tfrac{1}{2}-\kappa\bigr)\mathop{\mathrm{Re}}(f^{-1}\Theta(\alpha_{f}p_{f}-\lambda+q_{0})\bigr)\\ &\phantom{{}={}}{}+\tfrac{1}{2}\sigma\mathop{\mathrm{Re}}(f^{-1}\Theta(H-z)\bigr)-\tfrac{1}{2}\sigma f^{-1}\Theta(q_{1}+q_{2})-C_{5}Q\\ &\geq c_{2}\pi_{\mp}^{2}f^{-1+2\kappa}-\sigma\bigl(\tfrac{1}{2}-\kappa\bigr)\mathop{\mathrm{Re}}(f^{-1}\Theta(\alpha_{f}p_{f}-\lambda+q_{0})\bigr)-C_{6}Q.\end{split}$ (4.2)
We combine the fourth and the sixth terms as
$\begin{split}-\tfrac{1}{2}\sigma\Gamma\Theta(2\lambda\alpha_{f}-\alpha_{f}q_{0}-q_{0}\alpha_{f})\pm\tfrac{1}{2}\Gamma(2\lambda-q_{0}-\widetilde{q}_{0})\Theta&\geq\pm 2\Gamma\pi_{\mp}(\lambda-q_{0})\pi_{\mp}\Theta\\ &\geq 0.\end{split}$ (4.3)
As for the seventh term of (4.1), we bound it as
$\mp C_{4}\sigma\Gamma f^{-\rho}\Theta=\pm C_{4}\sigma\mathop{\mathrm{Im}}(f^{-\rho}\Theta(H-z))\mp C_{4}\sigma\mathop{\mathrm{Im}}(f^{-\rho}\Theta\alpha_{j}p_{j})\geq-C_{7}Q.$ (4.4)
We compute and bound the fifth term of (4.1) by using (1.1) and (4.4) as
$\begin{split}&\sigma\Gamma\mathop{\mathrm{Re}}\bigl(\alpha_{l}\alpha_{j}\Theta(\partial_{l}f)\ell_{jk}p_{k}\bigr)\\ &=\tfrac{1}{2}\sigma\Gamma\mathop{\mathrm{Re}}\bigl(\alpha_{l}\alpha_{j}\Theta(\partial_{l}f)\ell_{jk}p_{k}\bigr)+\tfrac{1}{2}\sigma\Gamma\mathop{\mathrm{Re}}\bigl(\alpha_{j}\alpha_{l}p_{k}\Theta(\partial_{l}f)\ell_{jk}\bigr)\\ &\geq\mp C_{8}\sigma\Gamma f^{-1}\Theta-C_{8}Q\\ &\geq-C_{9}Q.\end{split}$ (4.5)
Therefore by (4.1)–(4.5) we obtain
$\begin{split}&\sigma\mathop{\mathrm{Im}}\bigl((p_{f}\mp(z-q_{0}))^{*}\Theta(H-z)\bigr)\\ &\geq c_{2}\pi_{\mp}^{2}f^{-1+2\kappa}-\sigma\bigl(\tfrac{1}{2}-\kappa\bigr)\mathop{\mathrm{Re}}(f^{-1}\Theta(\alpha_{f}p_{f}-\lambda+q_{0})\bigr)-C_{9}Q.\end{split}$ (4.6)
Hence we are done. ∎
The second step for the proof of Proposition 4.1 is the following. We remark that Assumption 1.1 suffices for it.
Lemma 4.3.
Suppose Assumption 1.1, let $I\subset\mathbb{R}\setminus[m_{-},m_{+}]$ be a compact interval, and let $\kappa>0$. Then there exist $c,C>0$ and $a\in\mathbb{N}$ such that for any $z\in I_{\pm}$
$\begin{split}-\sigma\mathop{\mathrm{Re}}\bigl(f^{-1}\Theta(\alpha_{f}p_{f}-z+q_{0}+q_{1})\bigr)&\geq c(p_{f}-\alpha_{f}(z-q_{0}))^{*}f^{-1+2\kappa}(p_{f}-\alpha_{f}(z-q_{0}))\\ &\phantom{{}={}}{}-Cf^{-2-\rho+2\kappa}-C(H-z)^{*}f^{\rho+2\kappa}(H-z),\end{split}$
respectively.
Proof.
Fix a compact interval $I\subset\mathbb{R}\setminus[m_{-},m_{+}]$ and $\kappa>0$, and choose $a\in\mathbb{N}$ such that
$\sigma(\lambda-q_{0})\geq c_{1}>0\ \ \text{uniformly in }\lambda\in I\text{ and }|x|\geq 2^{a}.$
The below estimates are uniform in $z=\lambda+\mathrm{i}\Gamma\in I_{\pm}$. Set for short
$Q=f^{-2-\rho+2\kappa}+p_{j}f^{-2-\rho+2\kappa}p_{j}+(H-z)^{*}f^{\rho+2\kappa}(H-z).$
By Lemma 2.6 we see that $Q$ is absorbed into the last two terms of the asserted inequality, and thus we can regard it as a negligible error. In this proof we start with the first term on the right-hand side of the asserted inequality. Let us denote
$\Xi=\sigma(\lambda-q_{0})^{-1}f^{-1}\Theta,\quad\widetilde{\Xi}=\sigma(\lambda-\widetilde{q}_{0})^{-1}f^{-1}\Theta,$
and then by the Cauchy–Schwarz inequality and Assumption 1.1 we have
$(\alpha_{f}p_{f}-z+q_{0})^{*}f^{-1+2\kappa}(\alpha_{f}p_{f}-z+q_{0})$
$\leq C_{1}(\alpha_{f}p_{f}-z+q_{0})^{*}f^{-1}\Theta(\alpha_{f}p_{f}-z+q_{0})+C_{1}Q$
$=C_{1}(\alpha_{f}p_{f}-z+q)^{*}f^{-1}\Theta(\alpha_{f}p_{f}-z+q)$
$\phantom{{}={}}{}-2C_{1}\mathop{\mathrm{Re}}\bigl((q_{1}+q_{2})^{*}f^{-1}\Theta(\alpha_{f}p_{f}-z+q)\bigr)+C_{1}(q_{1}+q_{2})^{2}f^{-1}\Theta+C_{1}Q$
$\leq C_{2}(\alpha_{f}p_{f}-z+q)^{*}f^{-1}\Theta(\alpha_{f}p_{f}-z+q)+C_{2}Q$
$\leq C_{3}(\alpha_{f}p_{f}-z+q)^{*}\Xi(\alpha_{f}p_{f}-z+q)+C_{2}Q$
$=C_{3}\mathop{\mathrm{Re}}\bigl(p_{f}^{*}\alpha_{f}\Xi(\alpha_{f}p_{f}-z+q)\bigr)-C_{3}\mathop{\mathrm{Re}}\bigl((z-q_{0})^{*}\Xi(\alpha_{f}p_{f}-z+q_{0}+q_{1})\bigr)$
$\phantom{={}}{}+C_{3}\mathop{\mathrm{Re}}\bigl((q_{1}+q_{2})\Xi(\alpha_{f}p_{f}-z+q_{0}+q_{1})\bigr)-C_{3}\mathop{\mathrm{Re}}\bigl((z-q)^{*}\Xi q_{2}\bigr)+C_{2}Q$
$\leq C_{3}\mathop{\mathrm{Re}}\bigl(p_{f}^{*}\alpha_{f}\Xi(H-z)\bigr)-C_{3}\mathop{\mathrm{Re}}\bigl(p_{f}^{*}\alpha_{f}\Xi\alpha_{j}\ell_{jk}p_{k}\bigr)$
$\phantom{={}}{}-C_{3}\sigma\mathop{\mathrm{Re}}\bigl(f^{-1}\Theta(\alpha_{f}p_{f}-z+q_{0}+q_{1})\bigr)-C_{3}\Gamma\mathop{\mathrm{Im}}\bigl(\Xi(\alpha_{f}p_{f}-z+q_{0}+q_{1})\bigr)$
$\phantom{={}}{}+\tfrac{1}{4}(\alpha_{f}p_{f}-z+q_{0}+q_{1})^{*}f^{-1+2\kappa}(\alpha_{f}p_{f}-z+q_{0}+q_{1})+C_{4}Q$
$\leq-C_{3}\mathop{\mathrm{Re}}\bigl(p_{f}^{*}\alpha_{f}\Xi\alpha_{j}\ell_{jk}p_{k}\bigr)-C_{3}\sigma\mathop{\mathrm{Re}}\bigl(f^{-1}\Theta(\alpha_{f}p_{f}-z+q_{0}+q_{1})\bigr)$
$\phantom{={}}{}-C_{3}\Gamma\mathop{\mathrm{Im}}\bigl(\Xi(\alpha_{f}p_{f}-z+q_{0}+q_{1})\bigr)$
$\phantom{={}}{}+\tfrac{1}{2}(\alpha_{f}p_{f}-z+q_{0})^{*}f^{-1+2\kappa}(\alpha_{f}p_{f}-z+q_{0})+C_{5}Q.$
This implies that
$\begin{split}&(\alpha_{f}p_{f}-z+q_{0})^{*}f^{-1+2\kappa}(\alpha_{f}p_{f}-z+q_{0})\\ &\leq-2C_{3}\mathop{\mathrm{Re}}\bigl(p_{f}^{*}\alpha_{f}\Xi\alpha_{j}\ell_{jk}p_{k}\bigr)-2C_{3}\sigma\mathop{\mathrm{Re}}\bigl(f^{-1}\Theta(\alpha_{f}p_{f}-z+q_{0}+q_{1})\bigr)\\ &\phantom{={}}{}-2C_{3}\Gamma\mathop{\mathrm{Im}}\bigl(\Xi(\alpha_{f}p_{f}-z+q_{0}+q_{1})\bigr)+C_{6}Q.\end{split}$ (4.7)
We further compute and bound the right-hand side of (4.7). As for the first term, we can proceed, omitting $C_{3}$ and using (1.1), (2.4) and the Cauchy–Schwarz inequality, as
$\begin{split}&-2\mathop{\mathrm{Re}}\bigl(p_{f}^{*}\alpha_{f}\Xi\alpha_{j}\ell_{jk}p_{k}\bigr)\\ &=2p_{f}^{*}\alpha_{f}\Xi\alpha_{f}p_{f}-2\mathop{\mathrm{Re}}\bigl(p_{i}(\partial_{i}f)(\partial_{l}f)\alpha_{l}\Xi\alpha_{j}p_{j}\bigr)\\ &=2p_{f}^{*}\widetilde{\Xi}p_{f}-\mathop{\mathrm{Re}}\bigl(\alpha_{l}\alpha_{j}p_{i}(\partial_{i}f)(\partial_{l}f)\widetilde{\Xi}p_{j}\bigr)-\mathop{\mathrm{Re}}\bigl(\alpha_{j}\alpha_{l}p_{j}(\partial_{i}f)(\partial_{l}f)\widetilde{\Xi}p_{i}\bigr)\\ &=2p_{f}^{*}\widetilde{\Xi}p_{f}-\mathop{\mathrm{Re}}\bigl(\alpha_{l}\alpha_{j}p_{i}(\partial_{i}f)(\partial_{l}f)\widetilde{\Xi}p_{j}\bigr)-\mathop{\mathrm{Re}}\bigl(\alpha_{j}\alpha_{l}p_{i}(\partial_{i}f)(\partial_{l}f)\widetilde{\Xi}p_{j}\bigr)\\ &\phantom{={}}{}+\mathop{\mathrm{Im}}\bigl(\alpha_{j}\alpha_{l}(\partial_{i}(\partial_{i}f)(\partial_{l}f)\widetilde{\Xi})p_{j}\bigr)-\mathop{\mathrm{Im}}\bigl(\alpha_{j}\alpha_{l}(\partial_{j}(\partial_{i}f)(\partial_{l}f)\widetilde{\Xi})p_{i}\bigr)\\ &=\mathop{\mathrm{Im}}\bigl(\alpha_{j}\alpha_{l}(\Delta f)(\partial_{l}f)\widetilde{\Xi}p_{j}\bigr)+\mathop{\mathrm{Im}}\bigl(\alpha_{j}\alpha_{l}(\partial_{l}f)(\partial_{f}\widetilde{\Xi})p_{j}\bigr)\\ &\phantom{={}}{}-\mathop{\mathrm{Im}}\bigl(\alpha_{j}\alpha_{l}(\partial_{j}\partial_{i}f)(\partial_{l}f)\widetilde{\Xi}p_{i}\bigr)-\mathop{\mathrm{Im}}\bigl(\alpha_{j}\alpha_{l}(\partial_{i}f)(\partial_{j}\partial_{l}f)\widetilde{\Xi}p_{i}\bigr)\\ &\phantom{={}}{}-\mathop{\mathrm{Im}}\bigl(\alpha_{j}\alpha_{l}(\partial_{i}f)(\partial_{l}f)(\partial_{j}\widetilde{\Xi})p_{i}\bigr)\\ &=2\mathop{\mathrm{Im}}\bigl((\Delta f)\widetilde{\Xi}p_{f}\bigr)-\mathop{\mathrm{Im}}\bigl(\alpha_{f}(\Delta f)\Xi\alpha_{j}p_{j}\bigr)+2\mathop{\mathrm{Im}}((\partial_{f}\widetilde{\Xi})p_{f})\\ &\phantom{={}}{}-\mathop{\mathrm{Im}}\bigl(\alpha_{f}(\partial_{f}\Xi)\alpha_{j}p_{j}\bigr)+\mathop{\mathrm{Im}}\bigl(\alpha_{f}f^{-1}\Xi\alpha_{j}\ell_{ji}p_{i}\bigr)-\mathop{\mathrm{Im}}\bigl((\Delta f)\widetilde{\Xi}p_{f}\bigr)\\ &\phantom{={}}{}-2\mathop{\mathrm{Im}}((\partial_{f}\widetilde{\Xi})p_{f})+\mathop{\mathrm{Im}}(\alpha_{f}\alpha_{j}(\partial_{j}\widetilde{\Xi})p_{f})\\ &=-\mathop{\mathrm{Im}}\bigl(\alpha_{f}(f^{-1}-(\Delta f))\Xi(\alpha_{f}p_{f}-z+q_{0}+q_{1}+q_{2})\bigr)\\ &\phantom{={}}{}+\mathop{\mathrm{Im}}\bigl(\alpha_{f}(f^{-1}-(\Delta f))\Xi(H-z)\bigr)+\mathop{\mathrm{Im}}(\alpha_{f}\alpha_{j}\ell_{jk}(\partial_{k}\widetilde{\Xi})p_{f})\\ &\leq\tfrac{1}{4}C_{3}^{-1}(\alpha_{f}p_{f}-z+q_{0})^{*}f^{-1+2\kappa}(\alpha_{f}p_{f}-z+q_{0})+C_{7}Q.\end{split}$ (4.8)
On the other hand, the third term on the right-hand side of (4.7) is bounded by using the Cauchy–Schwarz inequality and (4.4) twice as
$\begin{split}&-2C_{3}\Gamma\mathop{\mathrm{Im}}\bigl(\Xi(\alpha_{f}p_{f}-z+q_{0}+q_{1})\bigr)\\ &\leq\tfrac{1}{8}(\alpha_{f}p_{f}-z+q_{0}+q_{1})^{*}f^{-1+2\kappa}(\alpha_{f}p_{f}-z+q_{0}+q_{1})+C_{8}\Gamma^{2}f^{-1+2\kappa}\\ &\leq\tfrac{1}{4}(\alpha_{f}p_{f}-z+q_{0})^{*}f^{-1+2\kappa}(\alpha_{f}p_{f}-z+q_{0})+C_{9}Q.\end{split}$ (4.9)
Therefore by (4.7), (4.8) and (4.9) we obtain the assertion. ∎
Proof of Proposition 4.1.
The assertion is clear from Lemmas 4.2 and 4.3. ∎
Proof of Theorem 1.9.
The assertion almost trivially follows from Proposition 4.1 and Theorem 1.5. We only remark that, in order to verify that all weighted norms finite, we use the boundedness of $R(z)$ as $L^{2}_{s}\to L^{2}_{s}$ for any $z\in\mathbb{C}\setminus\mathbb{R}$ and $s\in\mathbb{R}$. Thus we are done. ∎
4.2 Applications
At last we verify Corollaries 1.11, 1.12 and 1.13, as applications of Theorems 1.3, 1.5 and 1.9.
4.2.1 Proof of LAP
Proof of Corollary 1.11.
Take any $I$, $s$ and $\epsilon$ as in the assertion, and set $s^{\prime}=s-\epsilon$. Let us discuss on $I_{+}$ since the arguments for $I_{-}$ are completely the same. For any $a\in\mathbb{N}_{0}$ and $z,w\in I_{+}$ we decompose
$\begin{split}R(z)-R(w)&=\bigl(\chi_{a}R(z)\chi_{a}-\chi_{a}R(w)\chi_{a}\bigr)+\bigl(R(z)-\chi_{a}R(z)\chi_{a}\bigr)\\ &\phantom{{}={}}{}-\bigl(R(w)-\chi_{a}R(w)\chi_{a}\bigr),\end{split}$ (4.10)
see (2.7) for $\chi_{a}$, and $\bar{\chi}_{a}$ appearing below. By Theorem 1.5 we can bound the second term on the right-hand side of (4.10) uniformly in $a\in\mathbb{N}_{0}$ and $z\in I_{+}$ as
$\begin{split}\|R(z)-\chi_{a}R(z)\chi_{a}\|_{\mathcal{L}(L^{2}_{s},L^{2}_{-s})}&\leq\|f^{-s}R(z)\bar{\chi}_{a}f^{-s}\|_{\mathcal{L}(\mathcal{H})}\\ &\phantom{{}={}}{}+\|f^{-s}\bar{\chi}_{a}R(z)\chi_{a}f^{-s}\|_{\mathcal{L}(\mathcal{H})}\\ &\leq C_{1}2^{-a(s-s^{\prime})}\\ &=C_{1}2^{-a\epsilon}.\end{split}$ (4.11)
Similarly for the third term of (4.10), we have uniformly in $a\in\mathbb{N}_{0}$ and $w\in I_{+}$
$\|R(w)-\chi_{a}R(w)\chi_{a}\|_{\mathcal{L}(L^{2}_{s},L^{2}_{-s})}\leq C_{1}2^{-a\epsilon}.$ (4.12)
As for the first term of (4.10), we rewrite it as
$\chi_{a}R(z)\chi_{a}-\chi_{a}R(w)\chi_{a}$ $=\mathrm{i}\chi_{a}R(z)\alpha_{f}\chi_{a+1}^{\prime}R(w)\chi_{a}+(z-w)\chi_{a}R(z)\chi_{a+1}R(w)\chi_{a}$
$=\mathrm{i}\chi_{a}R(z)\pi_{f,+}\chi_{a+1}^{\prime}R(w)\chi_{a}-\mathrm{i}\chi_{a}R(z)\chi_{a+1}^{\prime}\pi_{f,-}R(w)\chi_{a}$
$\phantom{{}={}}{}+(z-w)\chi_{a}R(z)\pi_{f,+}\chi_{a+1}R(w)\chi_{a}$
$\phantom{{}={}}{}+(z-w)\chi_{a}R(z)\chi_{a+1}\pi_{f,-}R(w)\chi_{a}.$
Then by Theorems 1.5 and 1.9 uniformly in $a\in\mathbb{N}_{0}$ and $z,w\in I_{+}$
$\|\chi_{a}R(z)\chi_{a}-\chi_{a}R(w)\chi_{a}\|_{\mathcal{L}(L^{2}_{s},L^{2}_{-s})}\leq C_{2}2^{-a\epsilon}+C_{2}2^{a(1-\epsilon)}|z-w|.$ (4.13)
By (4.10), (4.11), (4.12) and (4.13), we obtain uniformly in $a\in\mathbb{N}_{0}$ and $z,w\in I_{+}$
$\|R(z)-R(w)\|_{\mathcal{L}(L^{2}_{s},L^{2}_{-s})}\leq C_{3}2^{-\epsilon n}+C_{2}2^{(1-\epsilon)n}|z-w|.$
Now, if $|z-w|\leq 1$, choose $a\in\mathbb{N}_{0}$ such that $2^{a}\leq|z-w|^{-1}\leq 2^{a+1}$, and then
$\|R(z)-R(w)\|_{\mathcal{L}(L^{2}_{s},L^{2}_{-s})}\leq C_{4}|z-w|^{\epsilon}.$
The same bound is trivial for |zw|>1|z-w|>1, and hence the Hölder continuity (1.14) is verified for $R(z)$. We can argue similarly for $p_{j}R(z)$.
The existence of the limits (1.15) follows from (1.14). In addition, by Theorem 1.5 $R_{\pm}(\lambda)$ and $p_{j}R(\lambda)$ map into $\mathcal{B}^{*}$, and then with the density argument they extend continuously as $\mathcal{B}\to\mathcal{B}^{*}$. Hence we are done. ∎
4.2.2 Proof of radiation condition bounds for limiting resolvents
Proof of Corollary 1.12.
The assertion is a direct consequence of Theorem 1.9 and Corollary 1.11. Hence we are done. ∎
4.2.3 Proof of Sommerfeld’s uniqueness
Proof of Corollary 1.13.
Let $\lambda\in\mathbb{R}\setminus[m_{-},m_{+}]$ and $\kappa\in(0,\rho/2)$. First, for any $\psi\in L^{2}_{1/2+\kappa}$ set $\phi=R_{\pm}(\lambda)\psi$. Then by Corollaries 1.11 and 1.12 we can see that $\phi\in\mathcal{B}^{*}$, and that the conditions 1 and 2 hold true. Conversely, let $\phi\in f^{\kappa}\mathcal{B}^{*}$ and $\psi\in f^{-\kappa}\mathcal{B}$, and assume the conditions 1 and 3. If we set
$\Phi=\phi-R_{\pm}(\lambda)\psi\in\psi\in f^{\kappa}\mathcal{B}^{*},$
then by Corollaries 1.11 and 1.12 we have
  1. 1.
    $(H-\lambda)\Phi=0$ in the distributional sense,
  2. 3.
    $\pi_{\mp}\Phi\in f^{-\kappa}\mathcal{B}_{0}^{*}$, or $(p_{f}\mp\sigma(\lambda-q_{0}))\Phi\in f^{-\kappa}\mathcal{B}_{0}^{*}$
Here we note that 1 and 3 imply $\pi_{\mp}\Phi\in\mathcal{B}_{0}^{*}$. To see this, it suffices to consider the case $(p_{f}\mp\sigma(\lambda-q_{0}))\Phi\in f^{-\kappa}\mathcal{B}_{0}^{*}$ in 3. For any $a\in\mathbb{N}$ let us compute
$\mathop{\mathrm{Re}}(\alpha_{f}\bar{\chi}_{a}^{\prime}(H-\lambda))$ $=\mathop{\mathrm{Re}}\bigl(\bar{\chi}_{a}^{\prime}(p_{f}\mp\sigma(\lambda-q_{0}))\bigr)\pm\sigma 2\mathop{\mathrm{Re}}\bigl(\pi_{\mp}\bar{\chi}_{a}^{\prime}(\lambda-q_{0})\bigr)$
$\phantom{{}={}}{}+\mathop{\mathrm{Re}}\bigl(\alpha_{f}\alpha_{j}\bar{\chi}_{a}^{\prime}\ell_{jk}p_{k}\bigr)+\mathop{\mathrm{Re}}\bigl(\alpha_{f}\bar{\chi}_{a}^{\prime}(q_{1}+q_{2})\bigr)$
$=\mathop{\mathrm{Re}}\bigl(\bar{\chi}_{a}^{\prime}(p_{f}\mp\sigma(\lambda-q_{0}))\bigr)\pm\sigma 2\pi_{\mp}\bar{\chi}_{a}^{\prime}(\lambda-q_{0})\pi_{\mp}$
$\phantom{{}={}}{}\mp\sigma\tfrac{1}{2}\bar{\chi}_{a}^{\prime}(q_{0}-\widetilde{q}_{0})+\mathop{\mathrm{Re}}\bigl(\alpha_{f}\bar{\chi}_{a}^{\prime}(q_{1}+q_{2})\bigr),$
and then, taking the expectation of the above identity on the state $\Phi$, we obtain as $a\to\infty$
$0$ $\leq c_{1}\langle\pi_{\mp}\Phi,\bar{\chi}_{a}^{\prime}\pi_{\mp}\Phi\rangle$
$\leq\bigl\langle\pi_{\mp}\Phi,\bar{\chi}_{a}^{\prime}(\lambda-q_{0})\pi_{\mp}\Phi\bigr\rangle$
$=\mp\sigma\tfrac{1}{2}\mathop{\mathrm{Re}}\bigl\langle\Phi,\chi_{a}^{\prime}(p_{f}\mp\sigma(\lambda-q_{0}))\Phi\bigr\rangle+\tfrac{1}{4}\bigl\langle\Phi,\chi_{a}^{\prime}(q_{0}-\widetilde{q}_{0})\Phi\bigr\rangle$
$\phantom{{}={}}{}\mp\sigma\tfrac{1}{2}\mathop{\mathrm{Re}}\bigl\langle\Phi,\alpha_{f}\chi_{a}^{\prime}(q_{1}+q_{2})\Phi\bigr\rangle$
$\to 0.$
Thus we obtain $\pi_{\mp}\Phi\in\mathcal{B}_{0}^{*}$. Finally let us further verify that $\Phi\in\mathcal{B}_{0}^{*}$. To see this, we compute for any $a\in\mathbb{N}$
$2\mathop{\mathrm{Im}}(\chi_{a}(H-\lambda))=\alpha_{f}\chi_{a}^{\prime}=\pm\chi_{a}^{\prime}\mp 2\pi_{\mp}\chi_{a}^{\prime},$
and this and 1 imply that as $a\to\infty$
$0\leq\langle\Phi,\bar{\chi}_{a}^{\prime}\Phi\rangle=2\langle\Phi,\bar{\chi}_{a}^{\prime}\pi_{\mp}\Phi\rangle\to 0.$
Thus we obtain $\Phi\in\mathcal{B}^{*}_{0}$, and then by Theorem 1.3 we conclude $\Phi=0$, or $\phi=R_{\pm}(\lambda)\psi$. Hence we are done. ∎
Acknowledgements
S.A. is supported by Forefront Physics and Mathematics Program to Drive Transformation (FoPM), a World-leading Innovative Graduate Study (WINGS) Program, the University of Tokyo. K.I. is partially supported by JSPS KAKENHI Grant Number 23K03163. A part of this work was done while K.I. visited Aarhus University. He would like to express his gratitude for kind hospitality by Arne Jensen and Erik Skibsted there.

### v2

@@
 The results of the paper would serve as a foundation for the stationary scattering theory of the Dirac operator.
+Keywords: Dirac operator, spectral theory, eigenfunctions, resolvent estimates
+Mathematics Subject Classification 2020: 35P10, 35Q41, 81Q15
 1 Settings and results
@@
 1.1.1 Introduction
-In this paper we investigate the Dirac equation in the $d$-dimensional configuration space $\mathbb{R}^{d}$ with $d\in\mathbb{N}=\{1,2,\ldots\}$. We write it in the form of evolution equation as
+In this paper we investigate the Dirac equation on the Euclidean space $\mathbb{R}^{d}$ of dimension $d\in\mathbb{N}=\{1,2,\ldots\}$. We write it in the form of evolution equation as
 $\partial_{t}\psi=-\mathrm{i}H\psi$
@@
   1. 1.
-    The function space $\mathcal{B}^{*}_{0}$ is optimal in the sense that there actually exists a non-trivial solution to (1.10) in $\mathcal{B}^{*}$, see Remark 1.14 for the construction.
-The largest function space so far is $L^{2}_{-1/2}(\mathbb{R}^{d})$ for $d=3$ due to Vogelsang [31], as far as the authors are aware of.
+    The function space $\mathcal{B}^{*}_{0}$ is optimal since we can in fact construct a non-trivial solution to (1.10) in $\mathcal{B}^{*}$ by using a WKB-type approximation and Corollary 1.13 below, see, e.g., [18, Proposition 4.18].
+So far the largest space where the generalized eigenfunctions are absent is $L^{2}_{-1/2}(\mathbb{R}^{d})$ for $d=3$ due to Vogelsang [31], as far as the authors are aware of.
   2. 2.
@@
 then $\phi=R_{\pm}(\lambda)\psi$, respectively.
-Remark 1.14.
-Using Corollary 1.13, we can construct a non-trivial solution to (1.10). In fact, for any $\psi\in L^{2}_{1/2+\kappa}$ with $\kappa\in(0,\rho/2)$ we set
-$\phi=R_{+}(\lambda)\psi-R_{-}(\lambda)\psi.$
-Then it satisfies (1.10) due to 1 of Corollary 1.13, and is non-trivial due to 2 of Corollary 1.13 and Theorem 1.3.
 2 Proof of Rellich’s theorem
-2.1 Reduction to three propositions
+2.1 Reduction to two propositions
 In this section we prove Theorem 1.3, following the scheme of Ito–Skibsted [16].
-We split the proof into three steps.
+We split the proof into two steps.
 Obviously, Theorem 1.3 is a consequence of the following propositions.
@@
 $(H-\lambda)\phi=0\ \ \text{in the distributional sense.}$
-Then there exists $\mu>0$ such that $\mathrm{e}^{\mu f}\phi\in\mathcal{B}_{0}^{*}$.
-Proposition 2.2.
-Let $\phi\in\mathcal{B}_{0}^{*}$ and $\lambda\in\mathbb{R}\setminus[m_{-},m_{+}]$ satisfy
-  1. 1.
-    $(H-\lambda)\phi=0$ in the distributional sense,
-  2. 2.
-    $\mathrm{e}^{\kappa f}\phi\in\mathcal{B}_{0}^{*}$ for some $\kappa>0$.
-Then for any $\mu>0$ one has $\mathrm{e}^{(\kappa+\mu)f}\phi\in\mathcal{B}^{*}_{0}$.
+Then for any $\kappa>0$ one has $\mathrm{e}^{\kappa f}\phi\in\mathcal{B}_{0}^{*}$.
 Proposition 2.2.
@@
 Then $\phi=0$ as a function on $\mathbb{R}^{d}$.
-We will prove these propositions in Sections 2.3, 2.4 and 2.5, respectively, after short preliminaries in Section $2.$
+We will prove these propositions in Sections 2.3 and 2.4, respectively, after short preliminaries in Section $2.$
 2.2 Preliminaries
@@
 Here we present formulas, to be repeatedly referred to in this and later sections, involving a general weight function.
-The proofs of Propositions 2.1, 2.2 and 2.3 depend on a commutator method sharing conjugate operators of the same form, however, with different weight functions.
+The proofs of Propositions 2.1 and 2.2 depend on a commutator method sharing conjugate operators of the same form, however, with different weight functions.
 Given a weight function $\Theta$, we introduce a radial differential operator
@@
 Note that the integral in (2.8) is a refinement of the Yosida approximation, satisfying
-$0<\int_{0}^{f}(1+\tau/2^{\nu})^{-1-\delta}\,\mathrm{d}\tau\leq\min\{f,\delta 2^{\nu}\}\ \ \text{for any }\nu\in\mathbb{N}_{0},$
+$0<\int_{0}^{f}(1+\tau/2^{\nu})^{-1-\delta}\,\mathrm{d}\tau\leq\min\{f,\delta^{-1}2^{\nu}\}\ \ \text{for any }\nu\in\mathbb{N}_{0},$
 and
@@
 $\begin{split}\theta^{\prime}=2\kappa+2\mu(1+f/2^{\nu})^{-1-\delta},\quad\theta^{\prime\prime}=-2(1+\delta)\mu 2^{-\nu}(1+f/2^{\nu})^{-2-\delta},\quad\ldots.\end{split}$
-In particular, noting $2^{-\nu}(1+f/2^{\nu})^{-1}\leq f^{-1}$, we can bound for any $k\in\mathbb{N}$
-$|(\theta-2\kappa f)^{(k)}|\leq C_{\delta,k}f^{-k}(\theta^{\prime}-2\kappa).$
+In particular, noting $2^{-\nu}(1+f/2^{\nu})^{-1}\leq f^{-1}$, we can bound for any $k=2,3,\ldots$
+$|(\theta-2\kappa f)^{\prime}|\leq C_{\delta,1}(\theta^{\prime}-2\kappa),\qquad|(\theta-2\kappa f)^{(k)}|\leq C_{\delta,k}f^{2-k}|\theta^{\prime\prime}|.$
 2.3 A priori exponential decay estimate
@@
 Lemma 2.7.
-Let $\lambda\gtrless m_{\pm}$, respectively, and let $\kappa=0$ in the definition (2.6) of $\Theta$.
-Then there exist $\mu,c,C>0$ and $a_{0}\in\mathbb{N}_{0}$ such that uniformly in $b>a\geq a_{0}$ and $\nu\in\mathbb{N}_{0}$
+Let $\lambda\gtrless m_{\pm}$, respectively, and fix any $\kappa_{0}\geq 0$.
+Then there exist $\mu,c,C>0$ and $a_{0}\in\mathbb{N}_{0}$ such that uniformly in $\kappa\in[0,\kappa_{0}]$, $b>a\geq a_{0}$ and $\nu\in\mathbb{N}_{0}$
 $\pm\mathop{\mathrm{Im}}\bigl(A_{\Theta}(H-\lambda)\bigr)$ $\geq c\min\{f^{-1},\theta^{\prime}\}\Theta-C\bigl(\chi_{a-1,a+1}^{2}+\chi_{b-1,b+1}^{2}\big)f^{-1}\mathrm{e}^{\theta}$
 $\phantom{={}}{}-(H-\lambda)\gamma(H-\lambda),$
-respectively, where $\gamma=\gamma_{a,b}$ is a certain function satisfying $\mathop{\mathrm{supp}}\gamma\subset\mathop{\mathrm{supp}}\chi_{a,b}$ and $|\gamma|\leq C_{a,b}$.
+respectively, where $\gamma=\gamma_{a,b}$ is a certain function, independent of $\kappa$ and $\nu$, satisfying $\mathop{\mathrm{supp}}\gamma\subset\mathop{\mathrm{supp}}\chi_{a,b}$ and $|\gamma|\leq C_{a,b}$.
 Remarks 2.8.
   1. 1.
-    We can see that the first term on the right-hand side is exactly from the first and second term of Lemma 2.4 with $\Theta$ of the form (2.6).
+    The first term on the right-hand side is from the first and second term of Lemma 2.3 with $\Theta$ of the form (2.6). It is exactly the minimum of the coefficients
   2. 2.
-    Let us outline the deduction of Proposition 2.1 from Lemma 2.8: Take the expectation of the above inequality on the eigenstate $\phi$, and the left-hand side and the last term on the right-hand side vanish; Let $b\to\infty$, and a contribution from $\chi_{b-1,b+1}^{2}f^{-1}\mathrm{e}^{\theta}$ vanishes by definition of $\mathcal{B}^{*}_{0}$; Then let $\nu\to\infty$, which replaces the Yosida approximation $\theta$ by $f$.
+    Let us outline the deduction of Proposition 2.1 from Lemma 2.7: Assume
+    $\kappa_{0}=\sup\{\kappa\geq 0;\ \mathrm{e}^{\kappa f}\phi\in\mathcal{B}^{*}_{0}\}<\infty,$ (2.9)
+    let $\mu>0$ be as above, and let $\kappa=0$ if $\kappa_{0}=0$, and $\kappa\in[0,\kappa_{0})$ with $\kappa+\mu>\kappa_{0}$ otherwise; Take the expectation of the above inequality on the eigenstate $\phi$, and the left-hand side and the last term on the right-hand side vanish; Let $b\to\infty$, and a contribution from $\chi_{b-1,b+1}^{2}f^{-1}\mathrm{e}^{\theta}$ vanishes by definition of $\mathcal{B}^{*}_{0}$; Let $\nu\to\infty$, which replaces the Yosida approximation $\theta$ by $f$; Then we obtain $\mathrm{e}^{(\kappa+\mu)f}\phi\in\mathcal{B}^{*}_{0}$, which contradicts the assumption.
 Proposition 2.1 thus follows.
+The second term on the right-hand side of Lemma 2.3 may be seen as a boundary contribution, as in Gauss’s divergence theorem.
 Proof.
 Let us only discuss the upper sign since the lower one can be done by the same manner.
-Fix $\lambda>m_{\pm}$ and $\kappa=0$ as in the assertion.
+Fix $\lambda>m_{+}$ and $\kappa_{0}\geq 0$ as in the assertion.
 For the moment we merely choose $a_{0}\in\mathbb{N}_{0}$ such that for some $c_{1}>0$
-$\lambda-q_{0}\geq c_{1}\ \ \text{uniformly in }\{x\in\mathbb{R}^{d};\ |x|\geq 2^{a_{0}}\}.$ (2.9)
-Then, except for the last part of the proof, the below estimates are all uniform in $\mu\in(0,1]$, $b>a\geq a_{0}$ and $\nu\in\mathbb{N}_{0}$, and thus until then $c_{*},C_{*}>0$ are independent of them.
+$\min\{\lambda-q_{0},\lambda-q_{0}-q_{1}\}\geq c_{1},\ \ \text{uniformly in }\{x\in\mathbb{R}^{d};\ |x|\geq 2^{a_{0}}\}.$ (2.10)
+Then, except for the last part of the proof, the below estimates are all uniform in $\kappa\in[0,\kappa_{0}]$, $\mu\in(0,1]$, $b>a\geq a_{0}$ and $\nu\in\mathbb{N}_{0}$, and thus until then $c_{*},C_{*}>0$ are independent of them.
 Only in the very last step we shall restrict ranges of all these parameters to verify the assertion, where we may also retake $a_{0}\in\mathbb{N}_{0}$ larger if necessary.
 In this proof we gather and absorb admissible error terms into
-$\begin{split}Q&=\bigl(\chi_{a,b}f^{-1}\theta^{\prime}+\chi_{a,b}f^{-1-\rho}+|\chi_{a,b}^{\prime}|\bigr)\mathrm{e}^{\theta}\\ &\phantom{{}={}}{}+p_{j}\bigl(\chi_{a,b}f^{-1}\theta^{\prime}+\chi_{a,b}f^{-1-\rho}+|\chi_{a,b}^{\prime}|\bigr)\mathrm{e}^{\theta}p_{j}+(H-\lambda)f^{1+\rho}\Theta(H-\lambda),\end{split}$ (2.10)
-which we prove at last is negligible.
+$\begin{split}Q&=\bigl(\chi_{a,b}|\theta^{\prime\prime}|+\chi_{a,b}f^{-1-\rho}+|\chi_{a,b}^{\prime}|\bigr)\mathrm{e}^{\theta}\\ &\phantom{{}={}}{}+p_{j}\bigl(\chi_{a,b}|\theta^{\prime\prime}|+\chi_{a,b}f^{-1-\rho}+|\chi_{a,b}^{\prime}|\bigr)\mathrm{e}^{\theta}p_{j}+(H-\lambda)f^{1+\rho}\Theta(H-\lambda).\end{split}$ (2.11)
+We will later prove that it is in fact negligible.
 Now let us start to compute the left-hand side of the asserted inequality. By Lemma 2.3 and the Cauchy–Schwarz inequality we first have
@@
 Recalling the expression (2.3) of $\ell_{jk}$, and using the Cauchy–Schwarz inequality and (2.10), we can combine and bound the first and second terms on the right-hand side of (2.12) as
-$\begin{split}&\mathop{\mathrm{Re}}\bigl(\alpha_{f}\theta^{\prime}\Theta p_{f}\bigr)+\mathop{\mathrm{Re}}\bigl(\alpha_{j}f^{-1}\Theta\ell_{jk}p_{k}\bigr)\\ &=\tfrac{1}{2}\mathop{\mathrm{Re}}\bigl(\alpha_{f}(\theta^{\prime}+f^{-1})\Theta p_{f}\bigr)-\tfrac{1}{2}\mathop{\mathrm{Re}}\bigl(\alpha_{f}(f^{-1}-\theta^{\prime})\Theta p_{f}\bigr)\\ &\phantom{{}={}}+\tfrac{1}{2}\mathop{\mathrm{Re}}\bigl(\alpha_{j}(\theta^{\prime}+f^{-1})\Theta\ell_{jk}p_{k}\bigr)+\tfrac{1}{2}\mathop{\mathrm{Re}}\bigl(\alpha_{j}(f^{-1}-\theta^{\prime})\Theta\ell_{jk}p_{k}\bigr)\\ &=\tfrac{1}{2}\mathop{\mathrm{Re}}\bigl(\alpha_{j}(\theta^{\prime}+f^{-1})\Theta p_{j}\bigr)-\tfrac{1}{2}\mathop{\mathrm{Re}}\bigl(\alpha_{j}(f^{-1}-\theta^{\prime})\Theta h_{jk}p_{k}\bigr)\\ &=\tfrac{1}{2}(\lambda-q)(\theta^{\prime}+f^{-1})\Theta+\tfrac{1}{2}\mathop{\mathrm{Re}}\bigl((\theta^{\prime}+f^{-1})\Theta(H-\lambda)\bigr)\\ &\phantom{{}={}}-\tfrac{1}{2}\mathop{\mathrm{Re}}\bigl(\mathrm{e}^{\theta/2}\chi_{a,b}(f^{-1}-\theta^{\prime})\alpha_{j}h_{jk}p_{k}\mathrm{e}^{\theta/2}\bigr)\\ &\geq\tfrac{1}{2}(\lambda-q_{0}-q_{1})(\theta^{\prime}+f^{-1})\Theta-\tfrac{1}{4}(\lambda-q_{0})|f^{-1}-\theta^{\prime}|\Theta\\ &\phantom{{}={}}-\tfrac{1}{4}\mathrm{e}^{\theta/2}p_{i}h_{il}\alpha_{l}\chi_{a,b}(\lambda-q_{0})^{-1}|f^{-1}-\theta^{\prime}|\alpha_{j}h_{jk}p_{k}\mathrm{e}^{\theta/2}-C_{2}Q,\end{split}$ (2.12)
+$\begin{split}&\mathop{\mathrm{Re}}\bigl(\alpha_{f}\theta^{\prime}\Theta p_{f}\bigr)+\mathop{\mathrm{Re}}\bigl(\alpha_{j}f^{-1}\Theta\ell_{jk}p_{k}\bigr)\\ &=\tfrac{1}{2}\mathop{\mathrm{Re}}\bigl(\alpha_{f}(\theta^{\prime}+f^{-1})\Theta p_{f}\bigr)-\tfrac{1}{2}\mathop{\mathrm{Re}}\bigl(\alpha_{f}(f^{-1}-\theta^{\prime})\Theta p_{f}\bigr)\\ &\phantom{{}={}}+\tfrac{1}{2}\mathop{\mathrm{Re}}\bigl(\alpha_{j}(\theta^{\prime}+f^{-1})\Theta\ell_{jk}p_{k}\bigr)+\tfrac{1}{2}\mathop{\mathrm{Re}}\bigl(\alpha_{j}(f^{-1}-\theta^{\prime})\Theta\ell_{jk}p_{k}\bigr)\\ &=\tfrac{1}{2}\mathop{\mathrm{Re}}\bigl(\alpha_{j}(\theta^{\prime}+f^{-1})\Theta p_{j}\bigr)-\tfrac{1}{2}\mathop{\mathrm{Re}}\bigl(\alpha_{j}(f^{-1}-\theta^{\prime})\Theta h_{jk}p_{k}\bigr)\\ &=\tfrac{1}{2}(\lambda-q)(\theta^{\prime}+f^{-1})\Theta+\tfrac{1}{2}\mathop{\mathrm{Re}}\bigl((\theta^{\prime}+f^{-1})\Theta(H-\lambda)\bigr)\\ &\phantom{{}={}}-\tfrac{1}{2}\mathop{\mathrm{Re}}\bigl(\chi_{a,b}(f^{-1}-\theta^{\prime})f^{2-d}\alpha_{j}h_{jk}p_{k}f^{d-2}\mathrm{e}^{\theta}\bigr)\\ &\geq\tfrac{1}{2}(\lambda-q_{0}-q_{1})(\theta^{\prime}+f^{-1})\Theta-\tfrac{1}{4}(\lambda-q_{0})|f^{-1}-\theta^{\prime}|\Theta\\ &\phantom{{}={}}-\tfrac{1}{4}\mathrm{e}^{\theta}f^{d-2}p_{i}h_{il}\alpha_{l}\chi_{a,b}(\lambda-q_{0})^{-1}|f^{-1}-\theta^{\prime}|f^{4-2d}\mathrm{e}^{-\theta}\alpha_{j}h_{jk}p_{k}f^{d-2}\mathrm{e}^{\theta}\\ &\phantom{{}={}}-C_{2}Q,\end{split}$ (2.13)
 where we have set
@@
 $\widetilde{\Xi}$ $=\chi_{a,b}(\lambda-\widetilde{q}_{0})^{-1}\bigl((f^{-1}-\theta^{\prime})^{2}+f^{-2-2\rho}\bigr)^{1/2},$
-cf. (1.6), and we proceed with the third term of (2.12) by using (1.1), (2.3) and the Cauchy–Schwarz inequality as
-$\begin{split}&-\tfrac{1}{4}\mathrm{e}^{\theta/2}p_{i}h_{il}\alpha_{l}\chi_{a,b}(\lambda-q_{0})^{-1}|f^{-1}-\theta^{\prime}|\alpha_{j}h_{jk}p_{k}\mathrm{e}^{\theta/2}\\ &\geq-\tfrac{1}{4}\mathrm{e}^{\theta/2}p_{i}h_{il}\alpha_{l}\Xi\alpha_{j}h_{jk}p_{k}\mathrm{e}^{\theta/2}\\ &=-\tfrac{1}{4}\mathop{\mathrm{Re}}\bigl(\alpha_{l}\alpha_{j}\widetilde{\Xi}\mathrm{e}^{\theta/2}h_{il}h_{jk}p_{i}p_{k}\mathrm{e}^{\theta/2}\bigr)-\tfrac{1}{4}\mathop{\mathrm{Im}}\bigl(\alpha_{l}\mathrm{e}^{\theta/2}(\partial_{i}h_{il}\Xi h_{jk})\alpha_{j}p_{k}\mathrm{e}^{\theta/2}\bigr)\\ &=-\tfrac{1}{4}\mathop{\mathrm{Re}}\bigl(\widetilde{\Xi}\mathrm{e}^{\theta/2}h_{ij}h_{jk}p_{i}p_{k}\mathrm{e}^{\theta/2}\bigr)-\tfrac{1}{4}\mathop{\mathrm{Im}}\bigl(\alpha_{l}(\partial_{i}h_{il}\Xi h_{jk})\alpha_{j}\mathrm{e}^{\theta}p_{k}\bigr)\\ &\phantom{{}={}}+\tfrac{1}{8}\mathop{\mathrm{Re}}\bigl(\alpha_{l}(\partial_{i}h_{il}\Xi h_{jk})\alpha_{j}(\partial_{k}f)\theta^{\prime}\mathrm{e}^{\theta}\bigr)\\ &\geq-\tfrac{1}{4}\mathop{\mathrm{Re}}\bigl(\widetilde{\Xi}\mathrm{e}^{\theta/2}\delta_{ik}p_{i}p_{k}\mathrm{e}^{\theta/2}\bigr)-C_{3}Q.\end{split}$ (2.14)
-Using (1.1) and the Cauchy–Schwarz inequality again, and noting $\alpha_{f}^{2}=1$, we continue (2.14) as
-$\begin{split}&-\tfrac{1}{4}\mathrm{e}^{\theta/2}p_{i}h_{il}\alpha_{l}\chi_{a,b}(\lambda-q_{0})^{-1}|f^{-1}-\theta^{\prime}|\alpha_{j}h_{jk}p_{k}\mathrm{e}^{\theta/2}\\ &\geq-\tfrac{1}{4}\mathop{\mathrm{Re}}\bigl(\widetilde{\Xi}\mathrm{e}^{\theta/2}\alpha_{i}\alpha_{k}p_{i}p_{k}\mathrm{e}^{\theta/2}\bigr)-C_{3}Q\\ &=-\tfrac{1}{4}\mathrm{e}^{\theta/2}p_{i}\alpha_{i}\Xi\alpha_{k}p_{k}\mathrm{e}^{\theta/2}+\tfrac{1}{4}\mathop{\mathrm{Im}}\bigl(\mathrm{e}^{\theta/2}\alpha_{i}(\partial_{i}\Xi)\alpha_{k}p_{k}\mathrm{e}^{\theta/2}\bigr)-C_{3}Q\\ &=-\tfrac{1}{4}\bigl(p_{i}\alpha_{i}+\tfrac{\mathrm{i}}{2}\alpha_{f}\theta^{\prime}\bigr)\Xi\mathrm{e}^{\theta}\bigl(\alpha_{k}p_{k}-\tfrac{\mathrm{i}}{2}\alpha_{f}\theta^{\prime}\bigr)+\tfrac{1}{4}\mathop{\mathrm{Im}}\bigl(\alpha_{i}(\partial_{i}\Xi)\mathrm{e}^{\theta}\alpha_{k}p_{k}\bigr)\\ &\phantom{{}={}}-\tfrac{1}{8}\mathop{\mathrm{Re}}\bigl(\alpha_{i}\alpha_{f}(\partial_{i}\Xi)\theta^{\prime}\mathrm{e}^{\theta}\bigr)-C_{3}Q\\ &\geq-\tfrac{1}{4}p_{i}\alpha_{i}\Xi\mathrm{e}^{\theta}\alpha_{k}p_{k}+\tfrac{1}{8}\alpha_{i}\alpha_{k}\bigl(\partial_{i}\Xi\theta^{\prime}(\partial_{k}f)\mathrm{e}^{\theta}\bigr)-\tfrac{1}{16}\Xi\theta^{\prime 2}\mathrm{e}^{\theta}-C_{4}Q\\ &\geq-\tfrac{1}{4}(\lambda-q+H-\lambda)\Xi\mathrm{e}^{\theta}(\lambda-q+H-\lambda)+\tfrac{1}{16}\Xi\theta^{\prime 2}\mathrm{e}^{\theta}-C_{5}Q\\ &\geq-\tfrac{1}{4}(\lambda-q)\Xi\mathrm{e}^{\theta}(\lambda-q)-C_{6}Q\\ &\geq-\tfrac{1}{4}(\lambda-q_{0}-2q_{1})|f^{-1}-\theta^{\prime}|\Theta-C_{7}Q.\end{split}$ (2.15)
-Hence by (2.11), (2.12) and (2.15) we obtain
-$\begin{split}\mathop{\mathrm{Im}}\bigl(A_{\Theta}(H-\lambda)\bigr)&\geq(c_{1}-q_{1})\min\{f^{-1},\theta^{\prime}\}\Theta-C_{8}Q.\end{split}$ (2.16)
-Next we discuss $Q$. Here we claim
-$\begin{split}Q&\leq C_{9}\bigl(f^{-1}\theta^{\prime}+f^{-1-\rho}\bigr)\Theta+C_{9}\bigl(\chi_{a-1,a+1}^{2}+\chi_{b-1,b+1}^{2}\big)f^{-1}\mathrm{e}^{\theta}\\ &\phantom{{}={}}{}+C_{9}(H-\lambda)f^{1+\rho}\Theta(H-\lambda),\end{split}$ (2.17)
-and for that it suffices to consider the second term of (2.10) since the contribution from $|\chi_{a,b}^{\prime}|\mathrm{e}^{\theta}$ is obviously absorbed into the second term of (2.17).
+cf. (1.6), and we proceed with the third term of (2.13) by using (2.14), (2.3) and $\alpha_{f}^{2}=1$ as
+$\begin{split}&-\tfrac{1}{4}\mathrm{e}^{\theta}f^{d-2}p_{i}h_{il}\alpha_{l}\chi_{a,b}(\lambda-q_{0})^{-1}|f^{-1}-\theta^{\prime}|f^{4-2d}\mathrm{e}^{-\theta}\alpha_{j}h_{jk}p_{k}f^{d-2}\mathrm{e}^{\theta}\\ &\geq-\tfrac{1}{4}\mathrm{e}^{\theta}f^{d-2}p_{i}h_{il}\alpha_{l}\Xi f^{4-2d}\mathrm{e}^{-\theta}\alpha_{j}h_{jk}p_{k}f^{d-2}\mathrm{e}^{\theta}\\ &=-\tfrac{1}{4}\mathrm{e}^{\theta/2}\bigl(p_{i}+\tfrac{\mathrm{i}}{2}(\partial_{i}f)\theta^{\prime}+\mathrm{i}(d-2)(\partial_{i}f)f^{-1}\bigr)h_{il}\alpha_{l}\Xi\alpha_{j}h_{jk}\\ &\phantom{{}={}}{}\cdot\bigl(p_{k}-\tfrac{\mathrm{i}}{2}(\partial_{k}f)\theta^{\prime}-\mathrm{i}(d-2)(\partial_{k}f)f^{-1}\bigr)\mathrm{e}^{\theta/2}\\ &=-\tfrac{1}{4}\mathrm{e}^{\theta/2}p_{i}h_{il}\alpha_{l}\Xi\alpha_{j}h_{jk}p_{k}\mathrm{e}^{\theta/2}\\ &\phantom{{}={}}{}+\tfrac{1}{4}\mathop{\mathrm{Im}}\bigl(\alpha_{f}\Xi(\theta^{\prime}+2(d-2)f^{-1})\mathrm{e}^{\theta/2}\alpha_{j}h_{jk}p_{k}\mathrm{e}^{\theta/2}\bigr)\\ &\phantom{{}={}}{}-\tfrac{1}{16}\widetilde{\Xi}(\theta^{\prime}+2(d-2)f^{-1})^{2}\mathrm{e}^{\theta}\\ &\geq-\tfrac{1}{4}\mathop{\mathrm{Re}}\bigl(\alpha_{l}\alpha_{j}\mathrm{e}^{\theta/2}\widetilde{\Xi}h_{il}h_{jk}p_{i}p_{k}\mathrm{e}^{\theta/2}\bigr)-\tfrac{1}{4}\mathop{\mathrm{Im}}\bigl(\alpha_{l}\alpha_{j}\mathrm{e}^{\theta/2}(\partial_{i}\widetilde{\Xi}h_{il}h_{jk})p_{k}\mathrm{e}^{\theta/2}\bigr)\\ &\phantom{{}={}}{}+\tfrac{1}{4}\mathop{\mathrm{Im}}\bigl(\widetilde{\Xi}(\theta^{\prime}+2(d-2)f^{-1})\mathrm{e}^{\theta/2}p_{f}\mathrm{e}^{\theta/2}\bigr)\\ &\phantom{{}={}}{}-\tfrac{1}{4}\mathop{\mathrm{Im}}\bigl(\alpha_{f}\Xi(\theta^{\prime}+2(d-2)f^{-1})\mathrm{e}^{\theta/2}\alpha_{j}\ell_{jk}p_{k}\mathrm{e}^{\theta/2}\bigr)\\ &\phantom{{}={}}{}-\tfrac{1}{16}\widetilde{\Xi}\theta^{\prime 2}\mathrm{e}^{\theta}-\tfrac{d-2}{4}\widetilde{\Xi}f^{-1}\theta^{\prime}\mathrm{e}^{\theta}-C_{3}Q.\end{split}$ (2.15)
+We continue to compute (2.15). Note that by (2.14), (2.3) and (2.4)
+$\partial_{i}h_{jk}=2f^{-1}\ell_{ij}(\partial_{k}f)+2f^{-1}(\partial_{j}f)\ell_{ik},\quad(\Delta f)=(d-1)f^{-1}.$ (2.16)
+Then, after some computations employing (1.1), Assumption 1.1, the Cauchy–Schwarz inequality, (2.16) and $\ell_{ij}\alpha_{i}\alpha_{j}=(d-1)$ along with (2.3), (2.4), (2.14) and $\alpha_{f}^{2}=1$ as above, we obtain
+$\begin{split}&-\tfrac{1}{4}\mathrm{e}^{\theta}f^{d-2}p_{i}h_{il}\alpha_{l}\chi_{a,b}(\lambda-q_{0})^{-1}|f^{-1}-\theta^{\prime}|f^{4-2d}\mathrm{e}^{-\theta}\alpha_{j}h_{jk}p_{k}f^{d-2}\mathrm{e}^{\theta}\\ &\geq-\tfrac{1}{4}\mathop{\mathrm{Re}}\bigl(\mathrm{e}^{\theta/2}\widetilde{\Xi}h_{ij}h_{jk}p_{i}p_{k}\mathrm{e}^{\theta/2}\bigr)-\tfrac{d-1}{2}\mathop{\mathrm{Im}}\bigl(\alpha_{f}\alpha_{j}\widetilde{\Xi}f^{-1}\mathrm{e}^{\theta/2}h_{jk}p_{k}\mathrm{e}^{\theta/2}\bigr)\\ &\phantom{{}={}}{}-\tfrac{1}{2}\mathop{\mathrm{Im}}\bigl(\alpha_{l}\alpha_{j}\widetilde{\Xi}f^{-1}\mathrm{e}^{\theta/2}h_{il}\ell_{ij}p_{f}\mathrm{e}^{\theta/2}\bigr)-\tfrac{1}{2}\mathop{\mathrm{Im}}\bigl(\alpha_{l}\alpha_{f}\widetilde{\Xi}f^{-1}\mathrm{e}^{\theta/2}h_{il}\ell_{ik}p_{k}\mathrm{e}^{\theta/2}\bigr)\\ &\phantom{{}={}}{}+\tfrac{d-1}{8}\widetilde{\Xi}f^{-1}\theta^{\prime}\mathrm{e}^{\theta}-\tfrac{1}{4}\mathop{\mathrm{Im}}\bigl(\alpha_{f}\Xi(\theta^{\prime}+2(d-2)f^{-1})\mathrm{e}^{\theta}\alpha_{j}\ell_{jk}p_{k}\bigr)\\ &\phantom{{}={}}{}-\tfrac{1}{16}\widetilde{\Xi}\theta^{\prime 2}\mathrm{e}^{\theta}-\tfrac{d-2}{4}\widetilde{\Xi}f^{-1}\theta^{\prime}\mathrm{e}^{\theta}-C_{4}Q\\ &\geq-\tfrac{1}{4}\mathop{\mathrm{Re}}\bigl(\mathrm{e}^{\theta/2}\widetilde{\Xi}\delta_{ik}p_{i}p_{k}\mathrm{e}^{\theta/2}\bigr)-\tfrac{1}{4}\mathop{\mathrm{Im}}\bigl(\alpha_{f}\Xi\theta^{\prime}\mathrm{e}^{\theta}\alpha_{j}\ell_{jk}p_{k}\bigr)\\ &\phantom{{}={}}{}-\tfrac{1}{16}\widetilde{\Xi}\theta^{\prime 2}\mathrm{e}^{\theta}-\tfrac{d-3}{8}\widetilde{\Xi}f^{-1}\theta^{\prime}\mathrm{e}^{\theta}-C_{5}Q.\end{split}$ (2.17)
+Let us compute the first term on the right-hand side of (2.17) as follows:
+$\begin{split}&-\tfrac{1}{4}\mathop{\mathrm{Re}}\bigl(\widetilde{\Xi}\mathrm{e}^{\theta/2}\delta_{ik}p_{i}p_{k}\mathrm{e}^{\theta/2}\bigr)\\ &=-\tfrac{1}{4}\mathop{\mathrm{Re}}\bigl(\widetilde{\Xi}\mathrm{e}^{\theta/2}\alpha_{i}\alpha_{k}p_{i}p_{k}\mathrm{e}^{\theta/2}\bigr)\\ &=-\tfrac{1}{4}\mathrm{e}^{\theta/2}p_{i}\alpha_{i}\Xi\alpha_{k}p_{k}\mathrm{e}^{\theta/2}+\tfrac{1}{4}\mathop{\mathrm{Im}}\bigl(\mathrm{e}^{\theta/2}\alpha_{i}(\partial_{i}\Xi)\alpha_{k}p_{k}\mathrm{e}^{\theta/2}\bigr)\\ &\geq-\tfrac{1}{4}\bigl(p_{i}\alpha_{i}+\tfrac{\mathrm{i}}{2}\alpha_{f}\theta^{\prime}\bigr)\Xi\mathrm{e}^{\theta}\bigl(\alpha_{k}p_{k}-\tfrac{\mathrm{i}}{2}\alpha_{f}\theta^{\prime}\bigr)-C_{6}Q\\ &=-\tfrac{1}{4}p_{i}\alpha_{i}\Xi\mathrm{e}^{\theta}\alpha_{k}p_{k}+\tfrac{1}{4}\mathop{\mathrm{Im}}\bigl(\alpha_{f}\Xi\theta^{\prime}\mathrm{e}^{\theta}\alpha_{k}p_{k}\bigr)-\tfrac{1}{16}\widetilde{\Xi}\theta^{\prime 2}\mathrm{e}^{\theta}-C_{6}Q\\ &=-\tfrac{1}{4}(\lambda-q+H-\lambda)\Xi\mathrm{e}^{\theta}(\lambda-q+H-\lambda)+\tfrac{1}{4}\mathop{\mathrm{Im}}\bigl(\alpha_{f}\Xi\theta^{\prime}\mathrm{e}^{\theta}\alpha_{f}p_{f}\bigr)\\ &\phantom{{}={}}+\tfrac{1}{4}\mathop{\mathrm{Im}}\bigl(\alpha_{f}\Xi\theta^{\prime}\mathrm{e}^{\theta}\alpha_{j}\ell_{jk}p_{k}\bigr)-\tfrac{1}{16}\widetilde{\Xi}\theta^{\prime 2}\mathrm{e}^{\theta}-C_{6}Q\\ &\geq-\tfrac{1}{4}(\lambda-q)\Xi\mathrm{e}^{\theta}(\lambda-q)+\tfrac{1}{8}\bigl(\partial_{i}(\partial_{i}f)\widetilde{\Xi}\theta^{\prime}\mathrm{e}^{\theta}\bigr)+\tfrac{1}{4}\mathop{\mathrm{Im}}\bigl(\alpha_{f}\Xi\theta^{\prime}\mathrm{e}^{\theta}\alpha_{j}\ell_{jk}p_{k}\bigr)\\ &\phantom{{}={}}-\tfrac{1}{16}\widetilde{\Xi}\theta^{\prime 2}\mathrm{e}^{\theta}-C_{7}Q\\ &\geq-\tfrac{1}{4}(\lambda-q_{0}-2q_{1})|f^{-1}-\theta^{\prime}|\Theta+\tfrac{1}{4}\mathop{\mathrm{Im}}\bigl(\alpha_{f}\Xi\theta^{\prime}\mathrm{e}^{\theta}\alpha_{j}\ell_{jk}p_{k}\bigr)\\ &\phantom{{}={}}+\tfrac{1}{16}\widetilde{\Xi}\theta^{\prime 2}\mathrm{e}^{\theta}+\tfrac{d-1}{8}\widetilde{\Xi}f^{-1}\theta^{\prime}\mathrm{e}^{\theta}-C_{8}Q.\end{split}$ (2.18)
+Hence by (2.12), (2.13), (2.17), (2.18), and (2.10) we obtain
+$\begin{split}\mathop{\mathrm{Im}}\bigl(A_{\Theta}(H-\lambda)\bigr)&\geq c_{2}\bigl(\min\{f^{-1},\theta^{\prime}\}+|f^{-1}-\theta^{\prime}|f^{-1}\theta^{\prime}\bigr)\Theta-C_{9}Q\\ &\geq c_{2}\min\{f^{-1},\theta^{\prime}\}\Theta-C_{9}Q.\end{split}$ (2.19)
+Note that we have dropped a positive contribution since it is not necessary here. However, it is essential in the proof of Proposition 2.2.
+Next we discuss $Q$. We claim that
+$\begin{split}Q&\leq C_{10}\bigl(|\theta^{\prime\prime}|+f^{-1-\rho}\bigr)\Theta+C_{10}\bigl(\chi_{a-1,a+1}^{2}+\chi_{b-1,b+1}^{2}\big)f^{-1}\mathrm{e}^{\theta}\\ &\phantom{{}={}}{}+C_{10}(H-\lambda)f^{1+\rho}\Theta(H-\lambda).\end{split}$ (2.20)
+For that it suffices to consider the second term of (2.11) since the contribution from $|\chi_{a,b}^{\prime}|\mathrm{e}^{\theta}$ is obviously absorbed into the second term of (2.20).
 However, the second term of (2.11) is obviously absorbed into the right-hand side of (2.20) due to Lemma 2.5 and the Cauchy–Schwarz inequality.
@@
 Therefore by (2.19) and (2.20) it follows that
-$\begin{split}\mathop{\mathrm{Im}}\bigl(A_{\Theta}(H-\lambda)\bigr)&\geq\bigl((c_{1}-q_{1})\min\{f^{-1},\theta^{\prime}\}-C_{10}f^{-1}\theta^{\prime}-C_{10}f^{-1-\rho}\bigr)\Theta\\ &\phantom{{}={}}{}-C_{10}\bigl(\chi_{a-1,a+1}^{2}+\chi_{b-1,b+1}^{2}\big)f^{-1}\mathrm{e}^{\theta}\\ &\phantom{{}={}}{}-C_{10}(H-\lambda)f^{1+\rho}\Theta(H-\lambda),\end{split}$
-and we finally discuss the coefficient of the first term on the right-hand side of (2.25).
-Now we restrict the ranges of parameters.
-We first retake $a_{0}\in\mathbb{N}_{0}$ larger if necessary, so that for some $c_{2}>0$
-$(c_{1}-q_{1})\geq 2c_{2}\ \ \text{uniformly for }|x|\geq 2^{a_{0}},$
-or on $\mathop{\mathrm{supp}}\Theta=\mathop{\mathrm{supp}}\chi_{a,b}$
-$(c_{1}-q_{1})\min\{f^{-1},\theta^{\prime}\}-C_{10}f^{-1}\theta^{\prime}-C_{10}f^{-1-\rho}$
-$\geq\min\bigl\{(c_{2}-C_{10}\theta^{\prime})f^{-1}+(c_{2}-C_{10}f^{-\rho})f^{-1},(c_{2}-C_{10}f^{-1})\theta^{\prime}+c_{2}\theta^{\prime}-C_{10}f^{-1-\rho}\bigr\}.$
-We further retake $a_{0}\in\mathbb{N}_{0}$ larger again if necessary, so that for some $c_{3}>0$
-$c_{2}-C_{10}f^{-\rho}\geq c_{3}\ \ \text{and}\ \ c_{2}-C_{10}f^{-1}\geq c_{3}$
-on the same subset. Then we take $\mu\in(0,1]$ small enough that for some $c_{4}>0$
-$c_{2}-C_{10}\theta^{\prime}>c_{4}.$
-Lastly, since we have for some $c_{5}>0$
-$c_{2}\theta^{\prime}-C_{10}f^{-1-\rho}\geq c_{5}\mu f^{-1-\delta}-C_{10}f^{-1-\rho},$
-letting $a_{0}\in\mathbb{N}_{0}$ larger again if necessary, we are done with the proof. ∎
+$\begin{split}\mathop{\mathrm{Im}}\bigl(A_{\Theta}(H-\lambda)\bigr)&\geq c_{2}\bigl(\min\{f^{-1},\theta^{\prime}\}-C_{11}|\theta^{\prime\prime}|-C_{11}f^{-1-\rho}\bigr)\Theta\\ &\phantom{{}={}}{}-C_{11}\bigl(\chi_{a-1,a+1}^{2}+\chi_{b-1,b+1}^{2}\big)f^{-1}\mathrm{e}^{\theta}\\ &\phantom{{}={}}{}-C_{11}(H-\lambda)f^{1+\rho}\Theta(H-\lambda),\end{split}$ (2.21)
+and finally it suffices to look at the coefficient of the first term on the right-hand side of (2.21).
+However, it is straightforward by restricting the ranges of parameters.
+We first retake $a_{0}\in\mathbb{N}_{0}$ larger if necessary, and then choose $\mu\in(0,1]$ sufficiently small.
+Then we can verify that on $\mathop{\mathrm{supp}}\Theta=\mathop{\mathrm{supp}}\chi_{a,b}$
+$\min\{f^{-1},\theta^{\prime}\}-C_{11}|\theta^{\prime\prime}|-C_{11}f^{-1-\rho}\geq c_{3}\min\{f^{-1},\theta^{\prime}\}.$
+Thus we are done with the proof. ∎
 Proof of Proposition 2.1.
-Let $\phi\in\mathcal{B}^{*}_{0}$ and $\lambda\in\mathbb{R}\setminus[m_{-},m_{+}]$ be as in the assertion, and let $\mu>0$ and $a_{0}\in\mathbb{N}_{0}$ be from Lemma 2.8.
+Let $\phi\in\mathcal{B}^{*}_{0}$ and $\lambda\in\mathbb{R}\setminus[m_{-},m_{+}]$ be as in the assertion, and we follow the strategy outlined in Remarks 2.8.
+Let $\kappa_{0}\in[0,\infty]$ be from (2.9), let $\mu>0$ and $a_{0}\in\mathbb{N}_{0}$ be from Lemma 2.7, and choose $\kappa\geq 0$ as in Remarks 2.8.
 Then, taking the expectation of the inequality from Lemma 2.7 on the state $\chi_{a-2,b+2}\phi$, we obtain that for $b>a\geq a_{0}$ and $\nu\in\mathbb{N}_{0}$
@@
 with $\bar{\chi}_{a}$ being from (2.7). Next we let $\nu\to\infty$ in (2.23), and then by Lebesgue’s monotone convergence theorem again it follows that
-$\bar{\chi}_{a}^{1/2}f^{-1/2}\mathrm{e}^{\mu f}\phi\in\mathcal{H}.$
-This implies $\mathrm{e}^{\mu f}\phi\in L^{2}_{-1/2}\subset\mathcal{B}^{*}_{0}$, and thus we are done. ∎
-2.4 A priori super-exponential decay estimate
-To prove Proposition 2.2 we next present to the following key lemma.
-Lemma 2.10.
-Let $\lambda\gtrless m_{\pm}$, respectively, and fix any $\kappa>0$ and $\mu>0$ in the definition (2.6) of $\Theta$.
-Then there exist $c,C>0$ and $a_{0}\in\mathbb{N}_{0}$ such that uniformly in $b>a\geq a_{0}$ and $\nu\in\mathbb{N}_{0}$
-$\pm\mathop{\mathrm{Im}}\bigl(A_{\Theta}(H-\lambda)\bigr)$ $\geq c\Theta-C\bigl(\chi_{a-1,a+1}^{2}+\chi_{b-1,b+1}^{2}\big)f^{-1}\mathrm{e}^{\theta}-(H-\lambda)\gamma(H-\lambda),$
-respectively, where $\gamma=\gamma_{a,b}$ is a certain function satisfying $\mathop{\mathrm{supp}}\gamma\subset\mathop{\mathrm{supp}}\chi_{a,b}$ and $|\gamma|\leq C_{a,b}$.
-Proof.
-The proof is very similar to that of Lemma 2.8, and we present only key steps, omitting details of the computations.
-Note that here the parameters are slightly different, and we squeeze positivity in a different manner from Lemma 2.8.
-We only discuss the upper sign.
-Fix any $\lambda>m_{+}$ and $\kappa,\mu>0$ as in the assertion.
-We first choose $a_{0}\in\mathbb{N}_{0}$ such that for some $c_{1}>0$
-$\lambda-q_{0}\geq c_{1}\ \ \text{and}\ \ \theta^{\prime}-f^{-1}\geq 2\kappa-f^{-1}\geq c_{1}$ (2.20)
-uniformly in $\{x\in\mathbb{R}^{d};\ |x|\geq 2^{a_{0}}\}$, but later we may retake it larger. All below estimates are uniform in $b>a\geq a_{0}$ and $\nu\in\mathbb{N}_{0}$.
-In this proof we set
-$\begin{split}Q&=\bigl(f^{-1}\chi_{a,b}+|\chi_{a,b}^{\prime}|\bigr)\mathrm{e}^{\theta}+p_{j}\bigl(f^{-1}\chi_{a,b}+|\chi_{a,b}^{\prime}|\bigr)\mathrm{e}^{\theta}p_{j}\\ &\phantom{{}={}}{}+(H-\lambda)f\Theta(H-\lambda),\end{split}$ (2.21)
-and now let us start to compute the left-hand side of the asserted inequality. Similarly to (2.11) and (2.12), here noting (2.20), we can deduce
-$\begin{split}\mathop{\mathrm{Im}}\bigl(A_{\Theta}(H-\lambda)\bigr)&\geq\tfrac{1}{2}(\lambda-q_{0}-q_{1})(\theta^{\prime}+f^{-1})\Theta-\tfrac{1}{4}(\lambda-q_{0})(\theta^{\prime}-f^{-1})\Theta\\ &\phantom{{}={}}-\tfrac{1}{4}\mathrm{e}^{\theta/2}p_{i}h_{il}\alpha_{l}\chi_{a,b}(\lambda-q_{0})^{-1}(\theta^{\prime}-f^{-1})\alpha_{j}h_{jk}p_{k}\mathrm{e}^{\theta/2}\\ &\phantom{{}={}}-C_{1}Q,\end{split}$ (2.22)
-with $h_{jk}$ defined as (2.13). To compute and bound the third term on the right-hand side of (2.22), we set
-$\Xi=\chi_{a,b}(\lambda-q_{0})^{-1}(\theta^{\prime}-f^{-1}),\quad\widetilde{\Xi}=\chi_{a,b}(\lambda-\widetilde{q}_{0})^{-1}(\theta^{\prime}-f^{-1}),$
-and then, similarly to (2.14), it follows that
-$\begin{split}&-\tfrac{1}{4}\mathrm{e}^{\theta/2}p_{i}h_{il}\alpha_{l}\chi_{a,b}(\lambda-q_{0})^{-1}(\theta^{\prime}-f^{-1})\alpha_{j}h_{jk}p_{k}\mathrm{e}^{\theta/2}\\ &\geq-\tfrac{1}{4}\mathop{\mathrm{Re}}\bigl(\widetilde{\Xi}\mathrm{e}^{\theta/2}\delta_{ik}p_{i}p_{k}\mathrm{e}^{\theta/2}\bigr)-C_{2}Q.\end{split}$ (2.23)
-We proceed similarly to (2.15), so that
-$\begin{split}&-\tfrac{1}{4}\mathrm{e}^{\theta/2}p_{i}h_{il}\alpha_{l}\chi_{a,b}(\lambda-q_{0})^{-1}(\theta^{\prime}-f^{-1})\alpha_{j}h_{jk}p_{k}\mathrm{e}^{\theta/2}\\ &\geq-\tfrac{1}{4}\bigl(p_{i}\alpha_{i}+\tfrac{\mathrm{i}}{2}\alpha_{f}\theta^{\prime}\bigr)\Xi\mathrm{e}^{\theta}\bigl(\alpha_{k}p_{k}-\tfrac{\mathrm{i}}{2}\alpha_{f}\theta^{\prime}\bigr)+\tfrac{1}{4}\mathop{\mathrm{Im}}\bigl(\alpha_{i}(\partial_{i}\Xi)\mathrm{e}^{\theta}\alpha_{k}p_{k}\bigr)\\ &\phantom{{}={}}-\tfrac{1}{8}\mathop{\mathrm{Re}}\bigl(\alpha_{i}\alpha_{f}(\partial_{i}\Xi)\theta^{\prime}\mathrm{e}^{\theta}\bigr)-C_{2}Q\\ &\geq\tfrac{1}{16}\Xi\theta^{\prime 2}\mathrm{e}^{\theta}-\tfrac{1}{4}p_{i}\alpha_{i}\Xi\mathrm{e}^{\theta}\alpha_{k}p_{k}-C_{3}Q\\ &\geq c_{2}\Theta-\tfrac{1}{4}(\lambda-q_{0}-2q_{1})(\theta^{\prime}-f^{-1})\Theta-C_{4}Q.\end{split}$ (2.24)
-Thus by (2.22) and (2.24) we obtain
-$\begin{split}\mathop{\mathrm{Im}}\bigl(A_{\Theta}(H-\lambda)\bigr)&\geq c_{2}\Theta+(c_{1}-q_{1})f^{-1}\Theta-C_{5}Q.\end{split}$ (2.25)
-We can control $Q$ as follows. In fact, similarly to (2.17), we can prove
-$\begin{split}Q&\leq C_{6}f^{-1}\Theta+C_{6}\bigl(\chi_{a-1,a+1}^{2}+\chi_{b-1,b+1}^{2}\big)f^{-1}\mathrm{e}^{\theta}\\ &\phantom{{}={}}{}+C_{6}(H-\lambda)f\Theta(H-\lambda),\end{split}$ (2.26)
-and hence by (2.25) and (2.26)
-$\begin{split}\mathop{\mathrm{Im}}\bigl(A_{\Theta}(H-\lambda)\bigr)&\geq(c_{2}-C_{7}f^{-1})f\Theta-C_{7}\bigl(\chi_{a-1,a+1}^{2}+\chi_{b-1,b+1}^{2}\big)f^{-1}\mathrm{e}^{\theta}\\ &\phantom{{}={}}{}-C_{7}(H-\lambda)f\Theta(H-\lambda).\end{split}$
-Finally by letting $a_{0}\in\mathbb{N}_{0}$ larger if necessary we are done with the proof. ∎
-Proof of Proposition 2.2.
-Fix $\phi\in\mathcal{B}^{*}_{0}$, $\lambda\in\mathbb{R}\setminus[m_{-},m_{+}]$ and $\kappa,\mu>0$ as in the assertion, and take $a_{0}\in\mathbb{N}_{0}$ as in Lemma 2.10.
-Taking the expectation of the inequality from Lemma 2.10 on the state $\chi_{a-2,b+2}\phi$, we obtain that for $b>a\geq a_{0}$ and $\nu\in\mathbb{N}_{0}$
-$\begin{split}\bigl\|\Theta^{1/2}\phi\bigr\|^{2}&\leq C_{a}\|\chi_{a-1,a+1}\phi\|^{2}+C_{\nu}2^{-\nu}\|\chi_{b-1,b+1}\mathrm{e}^{\theta}\phi\|^{2}.\end{split}$ (2.27)
-If we let $b\to\infty$ in (2.27), the second term on the right-hand side vanishes since $\mathrm{e}^{\kappa f}\phi\in\mathcal{B}^{*}_{0}$, and thus by Lebesgue’s monotone convergence theorem
-$\bigl\|(\bar{\chi}_{a}\mathrm{e}^{\theta})^{1/2}\phi\bigr\|^{2}$ $\leq C_{a}\|\chi_{a-1,a+1}\phi\|^{2}$ (2.28)
-with $\bar{\chi}_{a}$ being from (2.7). We next let $\nu\to\infty$ in (2.28), and then by Lebesgue’s monotone convergence theorem again
-$\bar{\chi}_{a}^{1/2}\mathrm{e}^{(\kappa+\mu)f}\phi\in\mathcal{H}.$
-This implies the assertion. ∎
+$\bar{\chi}_{a}^{1/2}f^{-1/2}\mathrm{e}^{(\kappa+\mu)f}\phi\in\mathcal{H}.$
+This implies $\mathrm{e}^{(\kappa+\mu)f}\phi\in L^{2}_{-1/2}\subset\mathcal{B}^{*}_{0}$, and thus a contradiction. We are done. ∎
 2.4 Absence of super-exponentially decaying eigenstates
-A key lemma for the proof of Proposition 2.3 is the following.
+To prove Proposition 2.2 we present to the following key lemma.
 Lemma 2.9.
@@
 Then there exist $c,C>0$ and $a_{0}\in\mathbb{N}_{0}$ such that uniformly in $\kappa\geq 1$ and $b>a\geq a_{0}$
-$\pm\mathop{\mathrm{Im}}\bigl(A_{\Theta}(H-\lambda)\bigr)$ $\geq c\kappa^{3}\Theta-C\kappa^{2}\bigl(\chi_{a-1,a+1}^{2}+\chi_{b-1,b+1}^{2}\big)f^{-1}\mathrm{e}^{2\kappa f}$
+$\pm\mathop{\mathrm{Im}}\bigl(A_{\Theta}(H-\lambda)\bigr)$ $\geq c\kappa^{2}f^{-1}\Theta-C\kappa^{2}\bigl(\chi_{a-1,a+1}^{2}+\chi_{b-1,b+1}^{2}\big)f^{-1}\mathrm{e}^{2\kappa f}$
 $\phantom{={}}{}-(H-\lambda)\gamma(H-\lambda),$
@@
 Proof.
-Again the proof is very similar to those of Lemmas 2.8 and 2.10, and we present only key steps.
-Only the upper sign is discussed.
-Fix any $\lambda>m_{+}$, and let $\mu=0$.
+The proof is very similar to that of Lemma 2.7, and we present only key steps, omitting details of the computations.
+Note that here we have to take particular care of dependence on the parameter $\kappa\geq 1$, and we squeeze positivity from a different term from Lemma 2.7.
+We only discuss the upper sign.
+Fix any $\lambda>m_{+}$ and $\mu=0$ as in the assertion.
 We first choose $a_{0}\in\mathbb{N}_{0}$ such that for some $c_{1}>0$
-$\lambda-q_{0}\geq c_{1}\ \ \text{and}\ \ \theta^{\prime}-f^{-1}=2\kappa-f^{-1}\geq c_{1}\kappa$ (2.29)
+$\min\{\lambda-q_{0},\lambda-q_{0}-q_{1}\}\geq c_{1}\ \ \text{and}\ \ \theta^{\prime}-f^{-1}=2\kappa-f^{-1}\geq c_{1}\kappa$ (2.24)
 uniformly in $\{x\in\mathbb{R}^{d};\ |x|\geq 2^{a_{0}}\}$ and $\kappa\geq 1$, but later we may retake it larger. All below estimates are uniform in $\kappa\geq 1$ and $b>a\geq a_{0}$.
-In this proof we set
-$\begin{split}Q&=\kappa^{2}\bigl(f^{-1}\chi_{a,b}+|\chi_{a,b}^{\prime}|\bigr)\mathrm{e}^{2\kappa f}+p_{j}\bigl(f^{-1}\chi_{a,b}+|\chi_{a,b}^{\prime}|\bigr)\mathrm{e}^{2\kappa f}p_{j}\\ &\phantom{{}={}}{}+\kappa(H-\lambda)f\Theta(H-\lambda),\end{split}$ (2.30)
-and then, similarly to (2.11) and (2.12), or (2.22), noting (2.29), we can deduce
-$\begin{split}\mathop{\mathrm{Im}}\bigl(A_{\Theta}(H-\lambda)\bigr)&\geq\tfrac{1}{2}(\lambda-q_{0}-q_{1})(2\kappa+f^{-1})\Theta-\tfrac{1}{4}(\lambda-q_{0})(2\kappa-f^{-1})\Theta\\ &\phantom{{}={}}-\tfrac{1}{4}\mathrm{e}^{\kappa f}p_{i}h_{il}\alpha_{l}\chi_{a,b}(\lambda-q_{0})^{-1}(2\kappa-f^{-1})\alpha_{j}h_{jk}p_{k}\mathrm{e}^{\kappa f}\\ &\phantom{{}={}}-C_{1}Q,\end{split}$ (2.31)
-with $h_{jk}$ being from (2.13). Set
-$\Xi=\chi_{a,b}(\lambda-q_{0})^{-1}(2\kappa-f^{-1}),\quad\widetilde{\Xi}=\chi_{a,b}(\lambda-\widetilde{q}_{0})^{-1}(2\kappa-f^{-1}),$
-and then, similarly to (2.14) or (2.23), the third term on the right-hand side of (2.31) is bounded as
-$\begin{split}&-\tfrac{1}{4}\mathrm{e}^{\kappa f}p_{i}h_{il}\alpha_{l}\chi_{a,b}(\lambda-q_{0})^{-1}(2\kappa-f^{-1})\alpha_{j}h_{jk}p_{k}\mathrm{e}^{\kappa f}\\ &\geq-\tfrac{1}{4}\mathop{\mathrm{Re}}\bigl(\widetilde{\Xi}\mathrm{e}^{\kappa f}\delta_{ik}p_{i}p_{k}\mathrm{e}^{\kappa f}\bigr)-C_{2}Q.\end{split}$
-We proceed similarly to (2.15) or (2.24) as
-$\begin{split}&-\tfrac{1}{4}\mathrm{e}^{\kappa f}p_{i}h_{il}\alpha_{l}\chi_{a,b}(\lambda-q_{0})^{-1}(2\kappa-f^{-1})\alpha_{j}h_{jk}p_{k}\mathrm{e}^{\kappa f}\\ &\geq-\tfrac{1}{4}\bigl(p_{i}\alpha_{i}+\mathrm{i}\kappa\alpha_{f}\bigr)\Xi\mathrm{e}^{2\kappa f}\bigl(\alpha_{k}p_{k}-\mathrm{i}\kappa\alpha_{f}\bigr)+\tfrac{1}{4}\mathop{\mathrm{Im}}\bigl(\alpha_{i}(\partial_{i}\Xi)\mathrm{e}^{2\kappa f}\alpha_{k}p_{k}\bigr)\\ &\phantom{{}={}}-\tfrac{1}{4}\kappa\mathop{\mathrm{Re}}\bigl(\alpha_{i}\alpha_{f}(\partial_{i}\Xi)\mathrm{e}^{2\kappa f}\bigr)-C_{2}Q\\ &\geq\tfrac{1}{4}\kappa^{2}\Xi\mathrm{e}^{2\kappa f}-\tfrac{1}{4}p_{i}\alpha_{i}\Xi\mathrm{e}^{2\kappa f}\alpha_{k}p_{k}-C_{3}Q\\ &\geq c_{2}\kappa^{3}\Theta-\tfrac{1}{4}(\lambda-q_{0}-2q_{1})(2\kappa-f^{-1})\Theta-C_{4}Q.\end{split}$ (2.32)
-Thus by (2.31) and (2.32)
-$\begin{split}\mathop{\mathrm{Im}}\bigl(A_{\Theta}(H-\lambda)\bigr)&\geq c_{2}\kappa^{3}\Theta+(c_{1}-q_{1})f^{-1}\Theta-C_{5}Q.\end{split}$ (2.33)
-Similarly to (2.17) or (2.26), we have
-$\begin{split}Q&\leq C_{6}\kappa^{2}f^{-1}\Theta+C_{6}\kappa^{2}\bigl(\chi_{a-1,a+1}^{2}+\chi_{b-1,b+1}^{2}\big)f^{-1}\mathrm{e}^{\theta}\\ &\phantom{{}={}}{}+C_{6}\kappa(H-\lambda)f\Theta(H-\lambda),\end{split}$ (2.34)
+Now let us start to compute the left-hand side of the asserted inequality. In this proof we set
+$\begin{split}Q&=\kappa^{2}\bigl(\chi_{a,b}f^{-1-\rho}+|\chi_{a,b}^{\prime}|\bigr)\mathrm{e}^{2\kappa f}+p_{j}\bigl(\chi_{a,b}f^{-1-\rho}+|\chi_{a,b}^{\prime}|\bigr)\mathrm{e}^{2\kappa f}p_{j}\\ &\phantom{{}={}}+\kappa(H-\lambda)f^{1+\rho}\Theta(H-\lambda).\end{split}$ (2.25)
+Then, similarly to (2.12) and (2.13), we can deduce, noting also (2.24),
+$\begin{split}\mathop{\mathrm{Im}}\bigl(A_{\Theta}(H-\lambda)\bigr)&\geq\tfrac{1}{2}(\lambda-q_{0}-q_{1})(2\kappa+f^{-1})\Theta-\tfrac{1}{4}(\lambda-q_{0})(2\kappa-f^{-1})\Theta\\ &\phantom{{}={}}-\tfrac{1}{4}\mathrm{e}^{2\kappa f}f^{d-2}p_{i}h_{il}\alpha_{l}\Xi f^{4-2d}\mathrm{e}^{-2\kappa f}\alpha_{j}h_{jk}p_{k}f^{d-2}\mathrm{e}^{2\kappa f}\\ &\phantom{{}={}}-C_{1}Q,\end{split}$ (2.26)
+where $h_{jk}$ is defined as (2.14), and
+$\Xi=\chi_{a,b}(\lambda-q_{0})^{-1}(2\kappa-f^{-1}),\quad\widetilde{\Xi}=\chi_{a,b}(\lambda-\widetilde{q}_{0})^{-1}(2\kappa-f^{-1}).$
+As for the third term on the right-hand side of (2.26), we repeat the arguments as in (2.15), (2.17) and (2.18), and then we obtain
+$\begin{split}&-\tfrac{1}{4}\mathrm{e}^{2\kappa f}f^{d-2}p_{i}h_{il}\alpha_{l}\widetilde{\Xi}f^{4-2d}\mathrm{e}^{-2\kappa f}\alpha_{j}h_{jk}p_{k}f^{d-2}\mathrm{e}^{2\kappa f}\\ &\geq-\tfrac{1}{4}(\lambda-q_{0}-2q_{1})(2\kappa-f^{-1})\Theta+\tfrac{\kappa}{2}\widetilde{\Xi}f^{-1}\mathrm{e}^{2\kappa f}-C_{3}Q.\end{split}$ (2.27)
+Hence by (2.26) and (2.27) it follows that
+$\begin{split}\mathop{\mathrm{Im}}\bigl(A_{\Theta}(H-\lambda)\bigr)&\geq c_{2}\kappa^{2}f^{-1}\Theta-C_{4}Q.\end{split}$ (2.28)
+Finally, similarly to (2.20), we can control $Q$ by Lemma 2.5 as
+$\begin{split}Q&\leq C_{5}\kappa^{2}f^{-1-\rho}\Theta+C_{5}\kappa^{2}\bigl(\chi_{a-1,a+1}^{2}+\chi_{b-1,b+1}^{2}\big)f^{-1}\mathrm{e}^{2\kappa f}\\ &\phantom{{}={}}{}+C_{5}\kappa(H-\lambda)f^{1+\rho}\Theta(H-\lambda),\end{split}$ (2.29)
 so that by (2.28) and (2.29)
-$\begin{split}\mathop{\mathrm{Im}}\bigl(A_{\Theta}(H-\lambda)\bigr)&\geq\kappa^{3}(c_{2}-C_{7}\kappa^{-1}f^{-1})f\Theta-C_{7}\kappa^{2}\bigl(\chi_{a-1,a+1}^{2}+\chi_{b-1,b+1}^{2}\big)f^{-1}\mathrm{e}^{\theta}\\ &\phantom{{}={}}{}-C_{7}\kappa(H-\lambda)f\Theta(H-\lambda).\end{split}$
-Letting $a_{0}\in\mathbb{N}_{0}$, we are done with the proof. ∎
+$\begin{split}\mathop{\mathrm{Im}}\bigl(A_{\Theta}(H-\lambda)\bigr)&\geq(c_{3}-C_{6}f^{-\rho})\kappa^{2}f^{-1}\Theta+C_{6}\kappa^{2}\bigl(\chi_{a-1,a+1}^{2}+\chi_{b-1,b+1}^{2}\big)f^{-1}\mathrm{e}^{2\kappa f}\\ &\phantom{{}={}}{}+C_{6}\kappa(H-\lambda)f^{1+\rho}\Theta(H-\lambda).\end{split}$
+Therefore, letting $a_{0}\in\mathbb{N}_{0}$ larger if necessary, we are done with the proof. ∎
 Proof of Proposition 2.2.
 Let $\phi\in\mathcal{B}^{*}_{0}$ and $\lambda\in\mathbb{R}\setminus[m_{-},m_{+}]$ be as in the assertion.
-Let $a_{0}\in\mathbb{N}_{0}$ be from Lemma 2.11, and take the expectation of the inequality from Lemma 2.11 on the state $\chi_{a-2,b+2}\phi$.
-Then for any $\kappa\geq 1$ and $b>a\geq a_{0}$
-$\kappa\|\chi_{a,b}^{1/2}\mathrm{e}^{\kappa f}\phi\|^{2}\leq C\|\chi_{a-1,a+1}\mathrm{e}^{\kappa f}\phi\|^{2}+C2^{-b}\|\chi_{b-1,b+1}\mathrm{e}^{\kappa f}\phi\|^{2}.$ (2.35)
-By the assumption the second term to the right of (2.35) vanishes as $b\to\infty$, and hence by Lebesgue’s monotone convergence theorem
-$\kappa\|\bar{\chi}_{a}^{1/2}\mathrm{e}^{\kappa f}\|^{2}$ $\leq C\|\chi_{b-1,b+1}\mathrm{e}^{\kappa f}\phi\|^{2},$
+Let $a_{0}\in\mathbb{N}_{0}$ be from Lemma 2.9, and taking the expectation of the inequality from Lemma 2.9 on the state $\chi_{a-2,b+2}\phi$.
+Then we have for any $\kappa\geq 1$ and $b>a\geq a_{0}$
+$\|\chi_{a,b}^{1/2}\mathrm{e}^{\kappa f}\phi\|^{2}\leq C\|\chi_{a-1,a+1}\mathrm{e}^{\kappa f}\phi\|^{2}+C2^{-b}\|\chi_{b-1,b+1}\mathrm{e}^{\kappa f}\phi\|^{2}.$ (2.30)
+If we let $b\to\infty$ in (2.30), the second term on the right-hand side vanishes since $\mathrm{e}^{\kappa f}\phi\in\mathcal{B}^{*}_{0}$, and thus by Lebesgue’s monotone convergence theorem
+$\bigl\|\bar{\chi}_{a}^{1/2}\mathrm{e}^{\kappa f}\phi\bigr\|^{2}\leq C\|\mathrm{e}^{\kappa f}\chi_{a-1,a+1}\phi\|^{2},$
 or
-$\kappa\|\bar{\chi}_{a}^{1/2}\mathrm{e}^{\kappa(f-2^{b+2})}\phi\|^{2}$ $\leq C\|\chi_{b-1,b+1}\phi\|^{2}.$ (2.36)
-Now we assume $\bar{\chi}_{b+2}\phi\neq 0$ as a function on $\mathbb{R}^{d}$.
-Then the left-hand side of (2.36) grows exponentially as $\kappa\to\infty$ whereas the right-hand side remains bounded.
-This is a contradiction, and thus $\bar{\chi}_{b+2}\phi=0$.
-Then by the unique continuation property [10] we conclude that $\phi=0$ as a function on $\mathbb{R}^{d}$.
+$\bigl\|\bar{\chi}_{a}^{1/2}\mathrm{e}^{\kappa(f-2^{a+2})}\phi\bigr\|^{2}\leq C\|\chi_{a-1,a+1}\phi\|^{2},$ (2.31)
+with $\bar{\chi}_{a}$ being from (2.7).
+Now we assume $\bar{\chi}_{a+2}\phi\not\equiv 0$ on $\mathbb{R}^{d}$.
+Then the left-hand side of (2.31) grows exponentially as $\kappa\to\infty$, whereas the right-hand side remains bounded.
+This is a contradiction, and thus $\bar{\chi}_{b+2}\phi\equiv 0$.
+Then by the unique continuation property [10] we conclude that $\phi\equiv 0$ on $\mathbb{R}^{d}$.
 We are done.
@@
 $\Xi=(\lambda-q_{0})^{-1}\bar{\chi}_{a}(f^{-1}\theta-\theta^{\prime}),\quad\widetilde{\Xi}=(\lambda-\widetilde{q}_{0})^{-1}\bar{\chi}_{a}(f^{-1}\theta-\theta^{\prime})$
-for short, and proceed similarly to (2.14) and (2.15) as
+for short, and proceed similarly to (2.15), (2.17) and (2.18) as
 $\begin{split}&-\tfrac{1}{4}p_{i}h_{il}\alpha_{l}(\lambda-q_{0})^{-1}\bar{\chi}_{a}(f^{-1}\theta-\theta^{\prime})\alpha_{j}h_{jk}p_{k}\\ &=-\tfrac{1}{4}\mathop{\mathrm{Re}}\bigl(\widetilde{\Xi}h_{ij}h_{jk}p_{i}p_{k}\bigr)-\tfrac{1}{4}\mathop{\mathrm{Im}}\bigl(\alpha_{l}(\partial_{i}h_{il}\Xi h_{jk})\alpha_{j}p_{k}\bigr)\\ &\geq-\tfrac{1}{4}\mathop{\mathrm{Re}}\bigl(\widetilde{\Xi}\delta_{ik}p_{i}p_{k}\bigr)-C_{3}Q\\ &=-\tfrac{1}{4}p_{i}\alpha_{i}\Xi\alpha_{k}p_{k}+\tfrac{1}{4}\mathop{\mathrm{Im}}\bigl(\alpha_{i}(\partial_{i}\Xi)\alpha_{k}p_{k}\bigr)-C_{3}Q\\ &\geq-\tfrac{1}{4}(\lambda-q_{0}-2q_{1})\bar{\chi}_{a}(f^{-1}\theta-\theta^{\prime})-C_{4}Q.\end{split}$ (3.5)
@@
 Therefore by (3.3), (3.6) and (3.7) it follows that
-$\mathop{\mathrm{Im}}\bigl(A_{\Theta}(H-z)\bigr)\geq(c_{1}-q_{1})\theta^{\prime}-\mathop{\mathrm{Re}}(\gamma(H-z))-C_{10}Q,$ (3.8)
+$\mathop{\mathrm{Im}}\bigl(A_{\Theta}(H-z)\bigr)\geq(c_{1}-q_{1})\theta^{\prime}-\mathop{\mathrm{Re}}(\gamma(H-z))-C_{9}Q,$ (3.8)
 Finally we can show by Lemmas 2.5 and 3.1 that
@@
 He would like to express his gratitude for kind hospitality by Arne Jensen and Erik Skibsted there.
+The authors would like to sincerely thank Kyohei Itakura for pointing out a serious error in the previous version of the paper.